### v1

# Equivalence of Privacy and Stability with Generalization Guarantees in Quantum Learning

Abstract
We present a unified information-theoretic framework to analyze the generalization performance of differentially private (DP) quantum learning algorithms. By leveraging the connection between privacy and algorithmic stability, we establish that $(\varepsilon,\delta)$-Quantum Differential Privacy (QDP) imposes a strong constraint on the mutual information between the training data and the algorithm’s output. We derive a rigorous, mechanism-agnostic upper bound on this mutual information for learning algorithms satisfying a 1-neighbor privacy constraint. Furthermore, we connect this stability guarantee to generalization, proving that the expected generalization error of any $(\varepsilon,\delta)$-QDP learning algorithm is bounded by the square root of the privacy-induced stability term. Finally, we extend our framework to the setting of an untrusted Data Processor, introducing the concept of Information-Theoretic Admissibility (ITA) to characterize the fundamental limits of privacy in scenarios where the learning map itself must remain oblivious to the specific dataset instance.
I Introduction
I-A Background and Motivation (Classical to Quantum)
A central goal of statistical learning theory is to understand when a hypothesis trained on finite data generalizes to unseen examples. In the classical domain, a profound equivalence exists between algorithmic stability—the insensitivity of a model to training set perturbations—and its generalization performance [3]. Concurrently, Differential Privacy (DP) [11] has emerged as the rigorous standard for stability in randomized algorithms, preventing overfitting even in adaptive data analysis [9, 1, 17].
As Quantum Machine Learning (QML) matures [2], it becomes urgent to establish classical-style guarantees—such as stability-to-generalization implications and privacy-based stability—for learning from quantum data [6, 19]. While Quantum DP (QDP) [23] successfully extends indistinguishability to quantum states, the intersection of QDP with statistical learning theory remains under-explored. The non-commutative nature of quantum information and the irreversibility of measurement necessitate a rigorous information-theoretic treatment.
I-B Our Framework and Main Contributions
In this work, we present a unified framework to determine the fundamental limits of generalization in quantum learning algorithms subject to privacy constraints. We adopt an information-theoretic approach, utilizing the mutual information between the training data and the algorithm’s output as the primary metric for stability [22].
To operationalize this study, we frame the learning protocol as an interaction between three distinct parties. The Respondent is the collection of individuals who contribute their private data to the training set; their primary concern is to prevent the leakage of their specific data instances. The Data Processor is the entity that aggregates this data and executes the quantum learning algorithm. Finally, the Investigator is the end-user who receives the algorithm’s output (the hypothesis) and whose goal is to minimize generalization error on unseen data. In the standard trusted setting, the Data Processor acts to protect data from the Investigator; in the untrusted setting, the Respondents require protection even from the Data Processor.
We structure our analysis around three pivotal logical steps: first establishing that stability governs generalization, then demonstrating that privacy enforces this stability, and finally extending these concepts to untrusted environments via Information-Theoretic Admissibility.
First step: Stability implies Generalizability. We begin by addressing the Investigator’s primary goal: to extract useful information that generalizes to unseen data. In the classical setting, the relationship between algorithmic stability and generalization has been extensively studied. Bousquet and Elisseeff [3] established a foundational connection, showing that if a learning algorithm is uniformly stable—meaning its output is insensitive to small perturbations in the training set—it guarantees tight bounds on the generalization error. This ensures that the empirical error observed on the training data is a reliable proxy for the true error on future data. Extending this classical intuition to the quantum domain, we prove that stable quantum algorithms possess similar strong generalization capabilities. By assuming a classical-quantum sub-Gaussian property for the loss function, we derive a bound on the expected generalization error in terms of the mutual information between the training data and the output. This result confirms that if an algorithm does not depend too heavily on any single quantum data point, it prevents overfitting.
Second step: Privacy implies Generalizability via Stability. Having established that stability is the key to generalization, we turn to the Respondent’s perspective: how do we ensure this stability in a rigorous manner? The answer lies in Differential Privacy. From the Respondent’s perspective, privacy necessitates that the Investigator cannot discern the inclusion or exclusion of any individual’s data. Differential privacy formalizes this requirement as a strong, mathematically precise form of stability. In the classical case, Dwork et al. [10] explicitly linked this to learning, demonstrating that differential privacy automatically ensures stability, and consequently, provides generalization guarantees. The main contribution of this manuscript is to generalize these results to the quantum learning scenario for the more general $(\varepsilon,\delta)$-DP setting. We propose a framework for a quantum learning algorithm which is "1-neighbor" $(\varepsilon,\delta)$-DP and study its stability by obtaining a rigorous upper-bound on the mutual information between the training data and the quantum output. By combining this with the result of the first step, we formally demonstrate that privacy is not merely a constraint, but a sufficient condition for generalization in quantum learning.
Third step: Dishonest Third Party and Information-Theoretic Admissibility. While the previous steps assume a trusted Data Processor who performs the learning algorithm, we also address the scenario where the third party, the Data Processor, cannot be trusted. In this untrusted Data Processor scenario, the learning algorithm must be independent of the specific training data instance to prevent the Data Processor from learning sensitive information. To rigorously characterize this setting, we introduce the concept of Information-Theoretic Admissibility (ITA). A learning algorithm is considered ITA with respect to a set of states if no other learning algorithm is strictly more informative—meaning no other algorithm allows the Investigator to recover the original information state via a CP-TP map.
If an algorithm is not ITA, the Investigator could potentially construct a "better" algorithm to extract more information, rendering the privacy definitions ineffective. We show that in the quantum setting, unlike the classical case where ITA implies full recoverability of information (meaning privacy is impossible under ITA), there exist non-trivial ITA algorithms where the security condition remains meaningful. Thus, the quantum setting provides a genuine advantage, validating the necessity of our privacy definitions even under the strict ITA condition.
I-C Organization of this Manuscript
The remainder of this manuscript is organized as follows. Section II establishes the requisite mathematical notation. In Section III, we present the general framework for quantum learning and formally define information-theoretic stability. Crucially, we prove that stability implies generalization (Step 2) by deriving an error bound based on a Classical-Quantum Sub-Gaussianity assumption. Section IV applies this framework to Differentially Private (DP) algorithms. We define the $1$-neighbor $(\varepsilon,\delta)$-DP condition and prove that privacy implies stability (Step 1), thereby establishing generalization guarantees for private quantum algorithms. We also compare our results with existing literature. Section V extends the framework to the untrusted Data Processor scenario, introducing Information-Theoretic Admissibility (ITA) to characterize privacy limits (Step 3). Finally, we conclude in Section VI.
II Notations
$\mathcal{D}(\mathcal{H})$ denotes the set of density operators on a finite-dimensional Hilbert space $\mathcal{H}$. The $\log$ function used in the discussions below is to the base $e$.
Method of Types.
Let $\mathcal{T}$ be a finite alphabet with $\left|\mathcal{T}\right|=d$. The type of a sequence $t\in\mathcal{T}^{n}$ is the frequency vector ${\bf f}=(f_{1},\ldots,f_{d})\in\mathbb{N}_{0}^{d}$ such that $\sum_{i}f_{i}=n$, where $f_{i}$ counts the occurrences of the $i$-th symbol of $\mathcal{T}$ in $t$. We denote the set of all such types by $T_{d}^{n}$. The type class $T_{\bf f}\subset\mathcal{T}^{n}$ is the set of all sequences sharing the same type $\bf f$.
[…]
III A General Framework for Quantum Learning Algorithms
III-A Learning Setup and Data Encoding
In this section, we establish a quantum learning framework motivated by [6] and [19], operationalized through the interaction between a Respondent (data contributor) and a Data Processor (algorithm executor), focusing on the non-private setting (i.e., without imposing any privacy constraint at this stage).
The process begins with the Respondent. The input consists of a classical dataset $s:=(z_{1},\cdots,z_{n})\in\mathcal{S}$ (where $n\in\mathbb{N},\left|\mathcal{Z}\right|>0$) collected from $n$ individual entries of the Respondent. Each Respondent contributes a corresponding quantum state $\rho_{z_{i}}\in\mathcal{D}(\mathcal{H}^{\bm{\mathfrak{Tr}}}\times\mathcal{H}^{\bm{\mathfrak{Te}}})$. Thus, the aggregate input state is $\rho_{s}:=\rho_{z_{1}}\otimes\cdots\otimes\rho_{z_{n}}$. This state consists of two parts: a training part over system $\bm{\mathfrak{Tr}}$ (accessible to the Data Processor) and a testing part over system $\bm{\mathfrak{Te}}$ (used for evaluation), which may be entangled. Each data point $z_{i}:=(x_{i},y_{i})$ maps an input $x_{i}$ to a label $y_{i}$ via an unknown concept function $f$.
The Data Processor receives this input, formalized as a classical-quantum state $\sum_{s}P_{S}(s)|s\rangle\langle s|\otimes\rho_{s}$, and executes a learning algorithm. We model this algorithm as a collection of quantum instruments [8] $\mathcal{N}:=\{\mathcal{N}^{(s)}:\mathcal{H}^{\bm{\mathfrak{Tr}}}\to\mathcal{H}^{B}\}_{s\in\mathcal{S}}$ which acts only on the input training system $\bm{\mathfrak{Tr}}$. The Processor produces an output on system $B$. Thus, the joint state between the input classical data system $S$, the input quantum testing system $\bm{\mathfrak{Te}}$, and the Data Processor’s output system $B$ is given by:
$\sigma^{S\bm{\mathfrak{Te}}B}_{\mathcal{N}}:=\sum_{s\in\mathcal{S}}P_{S}(s)|s\rangle\langle s|\otimes(\sigma^{\mathcal{N}}_{s})^{\textnormal{Te}B},$ (1)
where the output state conditioned on the Respondents’ input $s$ is,
$\sigma^{\mathcal{N}}_{s}:=\sum_{w\in\mathcal{W}}((\mathbb{I}^{\bm{\mathfrak{Te}}}\otimes\mathcal{N}^{(s)}_{w})(\rho_{s}))^{\bm{\mathfrak{Te}}B^{\prime}}\otimes|w\rangle\langle w|^{W}.$ (2)
Here, the Data Processor’s output system $B$ comprises a classical hypothesis system $W$ and a quantum residual system $B^{\prime}$, i.e., $B\equiv WB^{\prime}$. The action of the Data Processor’s instrument $\mathcal{N}^{(s)}$ is denoted as:
$\mathcal{N}^{(s)}(\rho_{s}):=\sum_{w\in\mathcal{W}}(\mathbb{I}^{\bm{\mathfrak{Te}}}\otimes\mathcal{N}^{(s)}_{w})(\rho_{s})\otimes|w\rangle\langle w|,$ (3)
where each $\mathcal{N}^{(s)}_{w}$ is a completely positive trace non-increasing map. This interaction is illustrated in Figure 1.
$\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$Respondent $\mathbf{\sum_{s}P_{S}(s)|s\rangle\langle s|\otimes\rho_{s}}$ $\mathbf{\{s,\rho_{s}\}}$$\mathbf{\mathcal{N}^{(s)}(\rho_{s})}$Learning AlgorithmData Processor
Figure 1: Privacy based learning framework.
III-B Stability of a Quantum Learning Algorithm
We now define stability for quantum learning algorithms. Intuitively, stability requires the learning outcome to remain invariant to single-entry modifications, thereby preventing the leakage of individual data points. Extending the classical information-theoretic framework of [22], which quantifies stability via mutual information, we formalize this notion below.
Definition 1.
(Stability) A quantum learning algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s}$ is defined to be $\gamma$-stable, if $\max_{P_{S}}I\left[S\bm{\mathfrak{Te}};WB^{\prime}\right]\leq\gamma$, where $I\left[S\bm{\mathfrak{Te}};WB^{\prime}\right]$ is calculated with respect to the classical-quantum state mentioned in (2).
The above definition provides a quantitative upper bound on the information that can be extracted from the Data Processor’s output $B$ ($B\equiv WB^{\prime}$) about the Respondent’s input dataset $S$ and $\bm{\mathfrak{Te}}$. Consequently, a small upper-bound implies that the algorithm’s output is not strongly dependent on any single training data point, indicating that the algorithm is information-theoretically stable.
III-C Stability Implies Generalizability For Quantum Learning Algorithms
In this section, we demonstrate that if the Data Processor employs a stable algorithm, the results generalize well to unseen data. For a quantum learning algorithm $\mathcal{N}=\{\mathcal{N}^{(s)}\}$, the joint state representing the Respondent’s input and the Data Processor’s output, mentioned in (1), can be expanded as:
$\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}:=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}P_{S}(s)\overset{S}{|s\rangle\langle s|}\otimes P^{\mathcal{N}}_{W\mid S}(w\mid s)\overset{W}{|w\rangle\langle w|}\otimes(\sigma^{\mathcal{N}}_{s,w})^{\bm{\mathfrak{Te}}B^{\prime}},$ (4)
where $P^{\mathcal{N}}_{W\mid S}(w\mid s):=\mathrm{Tr}\left[(\mathbb{I}^{\bm{\mathfrak{Te}}}\otimes\mathcal{N}^{(s)}_{w})(\rho_{s})\right]$ is the probability of the Data Processor selecting hypothesis $w$ given the dataset $s$, and $\sigma^{\mathcal{N}}_{s,w}$ is the normalized residual state $\frac{(\mathbb{I}^{\bm{\mathfrak{Te}}}\otimes\mathcal{N}^{(s)}_{w})(\rho_{s})}{P^{\mathcal{N}}_{W\mid S}(w\mid s)}$. In the following discussion, we define how to quantize the loss or error induced from the resultant state $\sigma^{\mathcal{N}}$.
In the earlier discussed quantum learning framework, the input data $s$ and output hypothesis $w$ induced by the quantum learning algorithm $\mathcal{N}$ are embedded into the output residue quantum state $\sigma^{\mathcal{N}}_{s,w}$. Therefore, to evaluate the performance of the Data Processor, we define the loss in terms of the expected value of observables with respect to the state $\sigma^{\mathcal{N}}$ produced by the Data Processor. In [6, 19], the authors consider a family of non-negative self-adjoint loss observables $\left\{L(s,w)\right\}_{\begin{subarray}{c}(w,s)\end{subarray}}$ which act on the quantum testing system $\bm{\mathfrak{Te}}$ and the output quantum system $B^{\prime}$. Using these loss observables, we define the following global loss operator,
$L^{S\bm{\mathfrak{Te}}WB^{\prime}}:=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}\overset{S}{|s\rangle\langle s|}\otimes\overset{W}{|w\rangle\langle w|}\otimes\overset{\bm{\mathfrak{Te}}B^{\prime}}{L(s,w)}.$ (5)
Based on the above description of the joint state $\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}$ and the loss operators $\{L(s,w)\}$, we now distinguish between the loss observed by the Data Processor on the training data (empirical) and the loss expected on unseen fresh data (true).
Definition 2 (Expected Empirical Loss [6, Definition $11$]).
The expected empirical loss $\hat{L}_{\rho}(\mathcal{N})$ captures the performance of the Data Processor’s algorithm on the dataset provided by the Respondent. It is the expectation over the joint distribution induced by the algorithm:
$\hat{L}_{\rho}(\mathcal{N})$ $:=\mathbb{E}_{(S,W)\sim{P}^{\mathcal{N}}_{SW}}[\mathrm{Tr}[L(S,W)(\sigma^{\mathcal{N}}_{S,W})^{\bm{\mathfrak{Te}}B^{\prime}}]]$
$=\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB^{\prime}}\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}].$
Definition 3 (Expected True Loss [19, Definition $19$]).
The expected true loss
$L_{\rho}(\mathcal{N})$ represents the generalization performance of the Data Processor’s algorithm. It evaluates the hypothesis $W$ generated by the Data Processor against a fresh dataset $\overline{S}$ independent of the training data $S$:
$L_{\rho}(\mathcal{N})$ $:=\mathbb{E}_{(\overline{S},\overline{W})\sim P_{S}\times{P}^{\mathcal{N}}_{{W}}}\left[\mathrm{Tr}\left[L(\overline{S},\overline{W})\left(\rho^{\bm{\mathfrak{Te}}}_{\overline{S}}\otimes(\sigma^{\mathcal{N}}_{\overline{W}})^{B^{\prime}}\right)\right]\right]$
$=\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB^{\prime}}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})].$
where for any $s$, we define $\rho^{\bm{\mathfrak{Te}}}_{s}:=\mathrm{Tr}_{\bm{\mathfrak{Tr}}}[\rho_{s}]$, for each $w$, we define $\sigma^{\mathcal{N}}_{w}:=\mathbb{E}_{S\sim P^{\mathcal{N}}_{S|W=w}}[\mathrm{Tr}_{\bm{\mathfrak{Te}}}[\sigma^{\mathcal{N}}_{S,w}]]$, and $\sigma^{S\bm{\mathfrak{Te}}}$ and $\sigma^{WB^{\prime}}_{\mathcal{N}}$ are the corresponding marginals of the state $\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}$ defined in (4).
Remark 1.
We adopt the definition of expected true loss as proposed in [19] and not that of [6]. The authors in [19] give a rigorous justification for Definition 3 and argue that the definition proposed by [6, Definition $12$] is not a correct definition for the expected true loss.
Based on these definitions, the expected generalization error is defined as the deviation between the Data Processor’s empirical performance and the true performance.
Definition 4 (Expected Generalization Error [19]).
The expected generalization error is:
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ $:=\left|\hat{L}_{\rho}(\mathcal{N})-L_{\rho}(\mathcal{N}))\right|$
$=\left|\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}]-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\right|.$
We will now bound $\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ in terms of $I\left[S\bm{\mathfrak{Te}};WB^{\prime}\right].$ To obtain such a bound in the classical setting [22] assumed that the loss function is sub-Gaussian. We will make a similar assumption for the loss operators $\{L(w,s)\}$ and the Data Processor’s output state. Towards this we make the following definition.
Definition 5 (Classical-Quantum Sub-Gaussianity).
For a fixed parameter $\alpha\in(0,\infty),$ the collection $\{L(w,s)\}$ of loss operators is said to be an $\alpha$ sub-Gaussian collection with respect to $\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}:=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}P_{S}(s)|s\rangle\langle s|\otimes P^{\mathcal{N}}_{W}(w\mid s)|w\rangle\langle w|\otimes\rho^{\bm{\mathfrak{Te}}}_{s}\otimes(\sigma^{\mathcal{N}}_{w})^{B^{\prime}},$ if for every $\lambda\in\mathbb{R},$ it satisfies,
$\mathbb{E}\left[\mathrm{Tr}\left[e^{\lambda\left(L(S,W)-\mathbb{E}\left[\mathrm{Tr}\left[L(S,W)\left(\rho^{\bm{\mathfrak{Te}}}_{S}\otimes(\sigma^{\mathcal{N}}_{W})^{B^{\prime}}\right)\right]\mathbb{I}^{\bm{\mathfrak{Te}}B^{\prime}}\right]\right)}\left(\rho^{\bm{\mathfrak{Te}}}_{S}\otimes(\sigma^{\mathcal{N}}_{W})^{B^{\prime}}\right)\right]\right]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2}},$ (6)
where the expectations are calculated with respect to the product distribution $P_{S}\times{P}^{\mathcal{N}}_{{W}}$. Note that (6) is equivalent to,
$\mathrm{Tr}\left[e^{\lambda\left(L^{S\bm{\mathfrak{Te}}WB}-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\mathbb{I}^{S\bm{\mathfrak{Te}}WB}\right)}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2}},$ (7)
where $L^{S\bm{\mathfrak{Te}}WB^{\prime}}$ is the global loss operator defined in (5).
Definition 5 naturally generalizes classical sub-Gaussianity. In the limit of trivial quantum systems ($\dim(\bm{\mathfrak{Te}})=\dim(B^{\prime})=1$), the operators $\sigma^{\mathcal{N}}_{W}$ and $\mathbb{I}^{B^{\prime}}$ become scalars, reducing $L(S,W)$ to a classical random loss function. Consequently, condition (6) collapses to the standard classical sub-Gaussian inequality
$\mathbb{E}_{(S,W)}[e^{\lambda(L(S,W)-\mathbb{E}[L(S,W)])}]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2}},$
with respect to $P_{S}\times{P}^{\mathcal{N}}_{{W}}$. We now present a theorem bounding the expected generalization error for quantum learning algorithms.
Theorem 1.
For a fixed $\alpha\in(0,\infty),$ if the loss operators for a quantum learning algorithm $\mathcal{N}$, satisfy Definition 5, then, we have,
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})\leq\sqrt{2\alpha^{2}I[S\bm{\mathfrak{Te}};WB^{\prime}]}.$ (8)
Proof.
See Appendix D for the proof.
Remark 2.
In [6, Theorem $17$], the authors impose the following two separate point-wise holding sub-Gaussianity requirements,
$\mathrm{Tr}\left[e^{\lambda(L(s,w)-f(s,w)\mathbb{I}^{\bm{\mathfrak{Te}}B^{\prime}})}(\rho^{\bm{\mathfrak{Te}}}_{s}\otimes(\sigma^{\mathcal{N}}_{s,w})^{B^{\prime}})\right]$ $\leq e^{\frac{\mu^{2}\lambda^{2}}{2}},\qquad\forall(s,w)\in\mathcal{S}\times\mathcal{W},$ (QMGF)
$\mathbb{E}_{S\sim P^{m}}\left[e^{\lambda(f(S,w)-\mathbb{E}_{\tilde{S}}[f(\tilde{S},w)])}\right]$ $\leq e^{\frac{\tau^{2}\lambda^{2}}{2}},\qquad\forall w\in\mathcal{W}.$ (CMGF)
for some fixed $\mu,\tau>0$, where $f(s,w)\coloneqq\mathrm{Tr}[L(s,w)(\rho^{\bm{\mathfrak{Te}}}_{s}\otimes(\sigma^{\mathcal{N}}_{s,w})^{B^{\prime}})]$. However, unlike Eqs. (QMGF) and (CMGF) in [6], Theorem 1 requires only a single condition mentioned in Definition 5. Crucially, our assumption holds in expectation over $P_{S}\times P_{W}$, rather than for worst-case pairs.
An important implication which follows from Theorem 1 and Definition 1 is the following corollary.
Corollary 1.
If the Data Processor’s learning algorithm $\mathcal{N}:=\{\mathcal{N}^{(s)}\}$ is $\gamma$-stable and the loss operators of $\mathcal{N}$ satisfy Definition 5, for a fixed $\alpha\in(0,\infty),$ then, its generalization error is upper bounded by $\sqrt{2\alpha^{2}\gamma}$.
This corollary establishes that if the Data Processor utilizes a stable algorithm, the generalization error is tightly controlled.
Remark 3.
Unlike the bounds in [6] and [19], our bound on $\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ relies on a single information-theoretic quantity. In scenarios where testing and training data are uncorrelated, the bound in [6] reduces to a purely classical term and therefore it will not account for the quantum system $B^{\prime}$. Therefore, from [6] it is not possible to show that stability implies generalizability. In contrast, Theorem 1 avoids this limitation, validating the definition of expected true loss proposed in [19] as its correct formulation. A justification for the correctness of (3) is also given in [19].
With this correct definition of true loss the bound obtained in [6, Corollary 23] translates to [19, Theorem 1]. In Appendix H, we make a comparison of Theorem 1 with [19, Theorem 1] for the case when $\alpha=\mu=\tau,$ where $\alpha,\mu$ and $\tau$ denote the sub-Gaussianity parameters appearing in (6), (QMGF), and (CMGF), respectively.
IV Generalization Guarantees for Differentially Private Quantum Learning
IV-A Trusted Setting and One-Neighbor $(\varepsilon,\delta)$-DP
This section examines the connection between privacy and generalization in quantum learning. Building on Section III, which linked information-theoretic stability to generalization, we demonstrate that differential privacy enforces this stability. Using the framework of Figure 2, we introduce the Investigator as the recipient of the output system $B$, generated by a Trusted Data Processor from the Respondent’s raw data ($S,\bm{\mathfrak{Te}},\bm{\mathfrak{Tr}}$). To prevent the reconstruction of individual entries, the Processor ensures the algorithm satisfies differential privacy, requiring output invariance under single-entry modifications. This constraint is mathematically equivalent to algorithmic stability (Definition 1), confirming privacy as a sufficient condition for generalization. We formalize this indistinguishability requirement below.
[…]
Figure 2: Privacy based learning framework.
Definition 6.
An algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$ is a $1$-neighbor $(\varepsilon,\delta)$-DP support consistent learning algorithm if it satisfies the following conditions:
  1. 1.
    Permutation Invariance: For all $s,s^{\prime}\in\mathcal{S}$ satisfying $T_{s}=T_{s^{\prime}}$, the algorithm satisfies the condition $\mathcal{N}^{(s)}(\rho_{s})=\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})$. This ensures that the algorithm’s output depends solely on the frequency of the data, not its specific ordering. This condition is natural in statistical learning, where the order of training examples is irrelevant to the hypothesis, and it further adds an extra layer of privacy.
  2. 2.
    Privacy: For every $s\overset{1}{\sim}s^{\prime}$ and $0\preceq\Lambda\preceq\mathbb{I}$, the following inequality holds:
    $\begin{split}\mathrm{Tr}[\Lambda\mathcal{N}^{(s)}(\rho_{s})]&\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})]+\delta,\\ \mathrm{Tr}[\Lambda\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})]&\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s)}(\rho_{s})]+\delta.\end{split}$ (9)
  3. 3.
    Support Consistency: For every $s\overset{1}{\sim}s^{\prime}$, the output supports are identical, i.e.,
    $\operatorname{supp}(\mathcal{N}^{(s)}(\rho_{s}))=\operatorname{supp}(\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})).$ (10)
    The support consistency condition (10) is not strictly necessary for all stability measures. See Remark 4 for more details.
Remark 4.
It is important to note that if stability is defined via the trace distance $\|\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}-(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\|_{1}$ rather than mutual information, (10) is not required, and similar bounds can be derived using our proof techniques. Note that for $\delta=0$, (10) is automatically satisfied.
IV-B Privacy Implies Stability: A Mutual-Information Bound
Definition 6 above implies that privacy guarantees extend to $k$-neighbors, albeit with degraded parameters.
Corollary 2.
If $\mathcal{N}$ satisfies Definition 6, then for any inputs $s\overset{k}{\sim}s^{\prime}$ ($k\geq 1$) and $0\preceq\Lambda\preceq\mathbb{I}$, we have, $\mathrm{Tr}[\Lambda\mathcal{N}^{(s)}(\rho_{s})]\leq e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})]+g_{k}(\varepsilon,\delta),$ where $g_{k}(\varepsilon,\delta):=\frac{e^{k\varepsilon}-1}{e^{\varepsilon}-1}\delta$ is assumed to be strictly less than $1$. The symmetric inequality holds by swapping $s$ and $s^{\prime}$.
Proof.
See Appendix E.
We now utilize the framework established in Section III to analyze the stability of quantum learning algorithms that satisfy Definition 6. For this analysis, we modify the framework by treating the quantum test data system $\bm{\mathfrak{Te}}$ as trivial (i.e., $\dim(\bm{\mathfrak{Te}})=1$).
Under this modification, the stability measure from Definition 1 simplifies to the mutual information between the training data $S$ and the output system $WB^{\prime}$. Therefore, in the theorem below, we derive an upper bound on $I[S;WB^{\prime}]$ for a quantum $(\varepsilon,\delta)$-differentially private (DP) learning algorithm. This derivation relies on the following assumption regarding the noise parameters $\varepsilon$ and $\delta$,
$g_{n(|\mathcal{Z}|-1)}(\varepsilon,\delta)<1,$ (11)
where for any $k\geq 1$, $g_{k}(\varepsilon,\delta)$ is defined in Corollary 2.
Theorem 2.
For $\varepsilon\in\left[\frac{1}{n},1)\right$., consider $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$ to be a $1$-neighbor $(\varepsilon,\delta)$-DP support consistent learning algorithm (see Definition 6) and satisfies the condition (11). Then, the following holds,
$I[S;WB^{\prime}]\leq(|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta),$ (12)
where, $n$ is the length of the training data and for some constant $m\in(0,1],$ $h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta):=\log\frac{1}{1-g_{{n(\left|\mathcal{Z}\right|-1)}}(\delta)}+\frac{2}{m}g_{n(\left|\mathcal{Z}\right|-1)}(\delta)$ and has a property that $h_{\left|\mathcal{Z}\right|}(\varepsilon,0)=0$.
Proof.
See Appendix F-A for the proof.
[…]
Corollary 3.
For $\varepsilon\in\left.[0,\frac{1}{n}\right)$, consider a learning algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$, which satisfies the properties mentioned in Definition 6 and the condition (11). Then,
$I[S;WB^{\prime}]\leq(|\mathcal{Z}|-1)\varepsilon n+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$
Proof.
See Appendix F-B for the proof.
Corollary 4.
[…]
$I[S;WB^{\prime}]\leq(|\mathcal{Z}|-1)\log\left(n+1\right).$
Proof.
See Appendix F-C for the proof.
Theorem 2 quantitatively links differential privacy to algorithmic stability by bounding the mutual information $I[S;WB^{\prime}]$ between the training data and the algorithm’s output. This bound is uniform and scales explicitly with dataset size $n$, alphabet size $|\mathcal{Z}|$, and privacy parameters $(\varepsilon,\delta)$. By translating $(\varepsilon,\delta)$-DP guarantees into a provable stability bound, the theorem establishes a direct connection between privacy and stability-based generalization controls.
Furthermore, Theorem 2 strictly generalizes [17, Proposition 2]: by taking a trivial quantum system ($\dim(B^{\prime})=1$) and setting $\delta=0$, the overhead $h_{\left|\mathcal{Z}\right|}$ vanishes, recovering the classical stability bound $(|\mathcal{Z}|-1)\log(ne\varepsilon)$.
Remark 5.
The upper-bound obtained in Theorem 2 is independent of $P_{S}$ and thus, Theorem 2 implies that if a quantum learning algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$ satisfies Definition 6, then $\mathcal{N}$ is $\left((|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+\right$. $\left.h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)\right)$-stable (see Definition 1). A similar observation also follows for Corollaries 3 and 4.
In Appendix B, we present a detailed comparison of Theorem 2 with [15, Proposition 10]. Additionally, in Appendix C, we provide an in-depth comparison of Theorem 2 with the results of [6, Appendix C.7] in the setting of untrusted Data Processors, a topic we will elaborate on in the subsequent section.
IV-C From Stability to Generalization: DP Generalization Guarantees
We now formally demonstrate that differential privacy guarantees generalization. By combining Theorem 2 with Theorem 1 (assuming a trivial system $\bm{\mathfrak{Te}}$), which bounds the expected generalization error via the square root of mutual information, we establish a direct link. Specifically, a $1$-neighbor $(\varepsilon,\delta)$-DP support consistent algorithm limits dependence on individual data points, thereby ensuring robust generalization. We formalize this result in the corollary below.
Corollary 5.
Consider a quantum learning algorithm $\mathcal{N}$ that is $1$-neighbor $(\varepsilon,\delta)$-DP support consistent (see Definition 6) with $\varepsilon\in[\frac{1}{n},1)$ and satisfies condition (11). If the loss operator satisfies (6) (or equivalently (7)) mentioned in Definition 5, for some $\alpha\in(0,\infty)$, then,
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})\leq\sqrt{2\alpha^{2}\mathcal{I}_{bound}},$
where $\mathcal{I}_{bound}=\left[(|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)\right]$ is the upper bound on the mutual information derived in Theorem 2.
V Untrusted Data Processor and Information Theoretic Admissibility (ITA)
V-A Untrusted Data Processor Model
In the previous sections, we assumed a trusted Data Processor model where the Data Processor reliably executes the privacy-preserving algorithm and releases only the privatized output. We now relax this assumption to address the Untrusted Data Processor scenario. Here, the Data Processor is considered adversarial and may attempt to leak or extract information about the training data $s$ beyond what is contained in the intended output system $B$. To address this privacy threat rigorously, we adopt a worst-case security model where the Data Processor and the Investigator collude or effectively act as a single adversarial entity.
In this setting, the Respondent does not grant the Data Processor access to the raw classical data $s$ directly. Instead, the Respondent provides access only to the set of encoded quantum states $\{\rho_{s}\}_{s}$. The Processor is tasked with running a learning algorithm to produce a hypothesis $w$. Since the Data Processor is untrusted, the learning algorithm is modeled as a single, fixed completely positive, trace-preserving (CP-TP) map $\mathcal{N}$ that must be independent of the specific input index $s$. The total state generated at the output of this process is, $\mathcal{N}(\rho_{s}):=\sum_{w\in\mathcal{W}}\mathcal{N}_{w}(\rho_{s})\otimes|w\rangle\langle w|,$ where each $\mathcal{N}_{w}$ is a completely positive trace non-increasing map summing to $\mathcal{N}$. To ensure privacy, the Respondent mandates that this map $\mathcal{N}$ must satisfy differential privacy constraints with respect to neighboring inputs. The formal definition of such a $1$-neighbor $(\varepsilon,\delta)$-DP support-consistent learning algorithm for an untrusted Data Processor is provided in Definition 8 of Appendix A.
V-B Information-Theoretic Admissibility (ITA): Motivation and Definition
While the Respondent prescribes a specific privacy-preserving algorithm $\mathcal{N}$, an adversarial Data Processor possessing the raw quantum inputs $\{\rho_{s}\}_{s}$ is not technically bound to execute $\mathcal{N}$. The Data Processor aims to extract the maximum possible information about $s$. Therefore, there exists a risk that the Data Processor might execute a strictly more informative algorithm $\mathcal{N}^{\prime}$ and then perform classical or quantum post-processing to simulate the output statistics of the prescribed algorithm $\mathcal{N}$, i.e., they can artificially stitch up the noise after performing a non-private learning algorithm.
If such a scenario is possible, the privacy guarantees calculated based on $\mathcal{N}$ (as mentioned in Definition 8) are rendered void, as the Data Processor effectively holds the information content of $\mathcal{N}^{\prime}$. To formalize this, we introduce the concept of ordering of informativeness between learning algorithms (Definition 9, Appendix A). We say that an algorithm $\mathcal{N}^{\prime}$ is more informative than $\mathcal{N}$ with respect to the ensemble $\{\rho_{s}\}_{s}$ if there exists a post-processing CP-TP map $\Gamma$ (a "simulation map") such that,
$\mathcal{N}(\rho_{s})=\Gamma\circ\mathcal{N}^{\prime}(\rho_{s}),\quad\forall s\in\mathcal{S}.$ (13)
If such a relation holds, the data-processing inequality [16] implies that the mutual information between the input and the output of $\mathcal{N}^{\prime}$ is strictly greater than or equal to that of $\mathcal{N}$.
To certify that a prescribed algorithm $\mathcal{N}$ dominates over every other algorithm $\mathcal{N}^{\prime}$ with respect to the collection $\{\rho_{s}\}_{s}$, we introduce the concept of Information-Theoretic Admissibility (ITA).
Definition 7 (Information-Theoretic Admissibility).
A learning algorithm $\mathcal{N}$ is ITA with respect to the set $\{\rho_{s}\}_{s}$ if there exists no other algorithm $\mathcal{N}^{\prime}$ that is strictly more informative than $\mathcal{N}$, i.e., there does not exist a $\Gamma$ which satisfies (13) for $\mathcal{N}^{\prime}$.
Essentially, if an algorithm is ITA, it implies that the Data Processor is already performing the optimal information extraction allowed by the quantum mechanics formalism.
V-C Quantum Advantage: Privacy under ITA
The imposition of the ITA condition reveals a fundamental divergence between classical and quantum privacy capabilities. We demonstrate that while classical ITA algorithms necessitate a total loss of privacy, quantum mechanics allows for algorithms that are both ITA and privacy-preserving.
V-C1 The Collapse of Classical Privacy:
In the classical domain, the encoded states $\rho_{s}$ effectively behave as probability distributions (or commutative states). In such a scenario, the lemma below shows that simultaneous ITA (optimality) and Differential Privacy (indistinguishability) are impossible.
Lemma 1.
Assume that all states $\{\rho_{s}\}_{s}$ commute (i.e., the classical setting). If there exists no reconstruction map $\Gamma$ such that $\Gamma\circ\mathcal{N}(\rho_{s})=\rho_{s}$, then the algorithm $\mathcal{N}$ is not ITA.
Proof.
See Appendix G
The implication of Lemma 1 is severe: classical ITA algorithms permit full reconstruction of the raw data. Since under ITA the Data Processor effectively holds the raw data, the output-based guarantees of Definition 8 are insufficient. Therefore, for classical ITA algorithms, the privacy condition in Definition 8 must be strengthened to require indistinguishability on the raw states $\rho_{s}$ directly, effectively substituting $\mathcal{N}(\rho_{s})$ with $\rho_{s}$.
V-C2 Quantum Privacy via Non-Commutativity:
In the quantum regime, the raw inputs $\{\rho_{s}\}_{s}$ may be non-commuting. Quantum mechanics dictates that non-orthogonal states cannot be perfectly distinguished, even by the optimal measurement. Therefore, an algorithm can be ITA—extracting the maximum physically permissible information—without yielding enough information to identify $s$ perfectly.
This allows for the existence of algorithms where the Data Processor performs the best possible measurement (ITA) but remains fundamentally limited by quantum uncertainty, thereby preserving privacy. We illustrate this with a concrete construction:
Example 3 (Quantum ITA algorithm).
Consider the states $\rho_{z}=|\phi_{z,p}\rangle\langle\phi_{z,p}|$ with $|\phi_{z,p}\rangle:=\sqrt{1-p}$ $|0\rangle+(-1)^{z}\sqrt{p}|1\rangle$. For a dataset $s=(z_{1},\dots,z_{n})$, define the encoded state $|e_{s}\rangle:=\bigotimes_{j=1}^{n}|\phi_{z_{j},1/2}\rangle$.Let $P_{k}$ be the projection onto the subspace spanned by $\{|e_{s}\rangle:|s|_{1}=k\}$ (i.e., the subspace of states corresponds to all strings $s$ with Hamming weight $k$). The untrusted Data Processor applies the projective measurement map $\{\mathcal{N}_{k}\}_{k}$ defined by $\mathcal{N}_{k}(\rho):=P_{k}\rho P_{k}$.
In Example 3, the proposed measurement strategy $\{\mathcal{N}_{k}\}_{k}$ satisfies the Information-Theoretic Admissibility (ITA) condition while maintaining the constraints for privacy. We analyze these properties below:
Justification for ITA: The projectors $\{P_{k}\}_{k}$ define the Hamming weight subspaces. In the orthogonal case ($p=1/2$), the input states $\{\rho_{s}\}_{s}$ are eigenstates of these projectors, making the measurement Quantum Non-Demolition (QND) [4] with $\mathcal{N}(\rho_{s})=\rho_{s}$. In the general case ($p\neq 1/2$), while the measurement may disturb the non-orthogonal states, it remains Information-Theoretically Admissible (ITA). This is because the Hamming weight measurement constitutes the optimal extraction of the ensemble’s geometric parameters; due to the permutation invariance of the encoded states, the Hamming weight acts as a sufficient statistic. Any attempt to extract strictly more information (e.g., individual bit positions) is physically precluded by qubit non-orthogonality, confirming the absence of any strictly superior map $\mathcal{N}^{\prime}$.
Justification for Privacy: The privacy guarantee stems from the distinction between optimal extraction (ITA) and perfect recovery. For $p\neq 1/2$, the qubit states $|\phi_{z,p}\rangle$ are non-orthogonal. Consequently, even though the algorithm extracts the maximal accessible information (the Hamming weight), the inherent quantum uncertainty defined by the Helstrom bound [14] prevents the Investigator from distinguishing the specific string $s$ within the projected subspace. This establishes a scenario where the algorithm is optimal (ITA) yet fundamentally privacy-preserving. Thus, it is meaningful to consider the security condition in Definition 8 as here in the above example, the ITA algorithm is inducing noise while performing the learning algorithm rather than artificially stitching it after.
V-D Distinction from Degradability:
Finally, it is crucial to distinguish ITA from the concept of quantum channel degradation [20, 5]. Degradability asks whether an algorithm can be simulated for any arbitrary input state. In contrast, ITA only asks whether the algorithm can be simulated on the specific training ensemble $\{\rho_{s}\}_{s}$. An algorithm might be non-degradable (secure in a general sense) but still simulable on the specific subspace occupied by the Respondent’s data. Therefore, privacy certification in the untrusted regime must be data-dependent, verifying admissibility explicitly against the geometry of the Respondent’s encoded states.
VI Conclusion
In this work, we established a rigorous information-theoretic framework linking Quantum Differential Privacy (QDP) to algorithmic stability and generalization. We derived a mechanism-agnostic upper bound on the Holevo information for $(\varepsilon,\delta)$-QDP algorithms, employing a grid-covering optimization to rigorously handle approximate privacy. This ensures stability guarantees that scale logarithmically with dataset size, recovering classical pure-DP results as a special case.
To bridge stability and generalization, we introduced the unified Classical-Quantum Sub-Gaussianity assumption. By capturing both classical sampling noise and intrinsic quantum measurement fluctuations in a single parameter, we derived compact bounds on the expected generalization error. Finally, via the concept of Information-Theoretic Admissibility (ITA), we demonstrated a fundamental quantum advantage: unlike the classical setting, quantum mechanics permits non-reversible, admissible algorithms that maintain strict privacy even against untrusted Data Processors.
Appendix A Technical Discussion related to Information-Theoretic Admissibility (ITA)
In the scenario where the Data Processor cannot be trusted, the privacy guarantee must hold against the Data Processor itself. Unlike the trusted setting where the Investigator only sees the classical output $w$, here the adversary has access to the quantum output $\mathcal{N}(\rho_{s})$. Consequently, the definition of differential privacy must be adapted to constrain the indistinguishability of the quantum states output by the learning map directly.
Definition 8.
An algorithm $\mathcal{N}$ is said to be a $1$-neighbor $(\varepsilon,\delta)$-DP support-consistent learning algorithm with an untrusted Data Processor if the following conditions hold.
  1. 1.
    Permutation Invariance: For all $s,s^{\prime}\in\mathcal{S}$ satisfying $T_{s}=T_{s^{\prime}}$, the algorithm satisfies the condition $\mathcal{N}(\rho_{s})=\mathcal{N}(\rho_{s^{\prime}})$.
  2. 2.
    Privacy: For every $s\overset{1}{\sim}s^{\prime}$ (see Section II) and $0\preceq\Lambda\preceq\mathbb{I}$, the following inequality holds:
    $\mathrm{Tr}[\Lambda\mathcal{N}(\rho_{s})]$ $\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}(\rho_{s^{\prime}})]+\delta,$
    $\mathrm{Tr}[\Lambda\mathcal{N}(\rho_{s^{\prime}})]$ $\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}(\rho_{s})]+\delta.$
  3. 3.
    Support Consistency: For every $s\overset{1}{\sim}s^{\prime}$, the output supports are identical, i.e.,
    $\operatorname{supp}(\mathcal{N}(\rho_{s}))=\operatorname{supp}(\mathcal{N}(\rho_{s^{\prime}})).$ (14)
To rigorously assess whether the privacy constraints are meaningful, we must ensure that the learning algorithm is not artificially “noisy” or suboptimal. We formalize this via the notion of admissibility. This concept allows us to order channels by their information content and identify when an algorithm extracts all accessible information from the input states.
Definition 9 (Information ordering of Algorithms).
Let $\mathcal{N}:=\{\mathcal{N}_{w}\}_{w}$ and $\mathcal{N}^{\prime}:=\{\mathcal{N}^{\prime}_{w}\}_{w}$ be two quantum learning algorithms and $\{\rho_{s}\}_{s}$ be a fixed set of input states to these algorithms. We say that $\mathcal{N}^{\prime}$ is more informative than $\mathcal{N}$ with respect to $\{\rho_{s}\}_{s}$ if there exists a family of CP-TP maps $\{\Gamma_{w}\}_{w}$ such that
$\Gamma_{w}\circ\mathcal{N}^{\prime}_{w}(\rho_{s})=\mathcal{N}_{w}(\rho_{s})\quad\forall\,(s,w)\in\mathcal{S}\times\mathcal{W}.$
Furthermore, $\mathcal{N}^{\prime}$ is said to be strictly more informative than $\mathcal{N}$ with respect to $\{\rho_{s}\}_{s}$ if $\mathcal{N}^{\prime}$ is more informative than $\mathcal{N}$, but the converse does not hold.
Appendix B Comparison between Theorem 2 and [15, Proposition 10]
In [15, Proposition 10], the authors derived an upper bound on the Holevo information for quantum $(\varepsilon,\delta)$-LDP quantum channels, as stated in Eq. (209) of [15]. However, one of the authors of [15] later clarified to the authors of the present paper [21] that the phrase “for quantum $(\varepsilon,\delta)$-LDP quantum channels” was a typographical error.
The corrected statement is as follows. If an algorithm $\mathcal{A}$ satisfies $\varepsilon$-QLDP, meaning that
$\mathrm{Tr}[M\mathcal{A}(\rho_{x})]\leq e^{\varepsilon}\mathrm{Tr}[M\mathcal{A}(\rho_{x^{\prime}})],\qquad\forall\,x,x^{\prime}\in\mathcal{X},\ \forall\,M:0\leq M\leq I,$ (15)
then, the following bound holds:
$I[X;B]_{\sigma}\leq\varepsilon\tanh\left(\frac{\varepsilon}{2}\right)=\varepsilon\left(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\right),$ (16)
where $I[X;B]_{\sigma}$ is the Holevo information computed with respect to the state $\sigma:=\sum_{x\in\mathcal{X}}P(x)|x\rangle\langle x|\otimes\mathcal{A}(\rho_{x})$. Here, their learning algorithm $\mathcal{A}$ can be considered as a map $\mathcal{N}$ given in (3), but the map does not depend on $s$. Hence, the reference [15] also adopts the Untrusted Data Processor scenario similar to the second part of [6].
In contrast, our main result, Theorem 2, provides an upper bound under a weaker assumption: the algorithm $\mathcal{A}$ satisfies $1$-neighbor $(\varepsilon,\delta)$-DP, i.e., the condition (9) holds for every $s\overset{1}{\sim}s^{\prime}$ (see Section II). When we set $\delta=0$ and $\mathcal{X}=\mathcal{S}$, the form of the constraint in (9) becomes identical to that in (15). However, our result applies this constraint only to $1$-neighboring pairs, whereas their result assumes it for all pairs $s\neq s^{\prime}\in\mathcal{S}$.
In fact, if we strengthen our assumption to match theirs—namely, require (9) for any distinct $s,s^{\prime}\in\mathcal{S}$—then part $(i)$ of [7, Corollary 3] recovers the same bound as (16), thereby aligning our result with the corrected version of [15, Proposition 10].
The proof of Theorem 1 formally establishes the connection between algorithmic stability and generalizability by treating the mutual information $I[S;WB^{\prime}]$ as a proxy for stability. We first demonstrate that the expected generalization error is fundamentally limited by the square root of the information the algorithm leaks about the training data $S$, i.e., we have the following,
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ $\leq\sqrt{2\alpha^{2}I[S;WB^{\prime}]}.$ (17)
Here, $I[S;WB^{\prime}]$ quantifies the dependence of the output hypothesis on the specific training set; a lower value implies that the algorithm is "stable" and not overfitting to individual data points. The crucial link to Theorem 2 is that it provides the explicit upper bound on this stability measure derived solely from the privacy constraints. Finally, by substituting the bound from Theorem 2 into (17), we mathematically confirm that the rigorous stability imposed by $(\varepsilon,\delta)$-DP directly suppresses the generalization error, by preventing the algorithm from depending too heavily on any single data point and ensuring that the learned hypothesis performs well on unseen data.
Appendix C Comparison between Theorem 2 and [6, Appendix C.7]
The reference [6] studies the local differential privacy of learning algorithms in two settings. That is, their discussion is composed of two parts, the first part starting with “First” and the second part starting with “Next”.
C-A Assumption in [6, Appendix C.7]
Their first part discusses the Holevo information under a certain condition. However, a careful examination of the proof in [6, Appendix C.7] reveals that the argument relies on a stronger assumption than their statement as follows. In this place, the authors claim to prove the following bound on the Holevo information.
$I(\text{test};\text{hyp})_{\sigma^{\mathcal{A}}_{(s,w)}}\leq 2\varepsilon(1-e^{-\varepsilon})\sqrt{2I(\text{test};\text{train})_{\rho^{\mathcal{A}}_{(s,w)}}},$ (18)
under the assumption that the channel $\Lambda^{\mathcal{A}}_{s,w}:\mathcal{H}^{\mbox{train}}\to\mathcal{H}^{\mbox{hyp}}$ is $\varepsilon$-LDP, i.e.,
$\mathrm{Tr}\left[M\Lambda^{\mathcal{A}}_{s,w}(\rho^{\mbox{train}}_{1})\right]\leq e^{\varepsilon}\mathrm{Tr}\left[M\Lambda^{\mathcal{A}}_{s,w}(\rho^{\mbox{train}}_{2})\right],$ (19)
for all $0\preceq M\preceq\mathbb{I}^{\mbox{hyp}}$ and $\rho^{\mbox{train}}_{1},\rho^{\mbox{train}}_{2}\in\mathcal{D}(\mathcal{H}^{\mbox{train}})$.
However, a closer examination of their proof reveals that the argument implicitly depends on a stronger condition, namely
$\mathrm{Tr}\left[O\big(\mathbb{I}^{\mathrm{test}}\otimes\Lambda^{\mathcal{A}}_{s,w}\big)\big(\rho^{\mathrm{test;train}}_{1}\big)\right]\leq e^{\varepsilon},\mathrm{Tr}\left[O\big(\mathbb{I}^{\mathrm{test}}\otimes\Lambda^{\mathcal{A}}_{s,w}\big)\big(\rho^{\mathrm{test;train}}_{2}\big)\right],$ (20)
for all $0\preceq O\preceq\mathbb{I}^{\mathrm{test;hyp}}$ and $\rho^{\mathrm{test;train}}_{1},\rho^{\mathrm{test;train}}_{2}\in\mathcal{D}(\mathcal{H}^{\mathrm{test;train}})$. In other words, the proof appears to require that $\Lambda^{\mathcal{A}}_{s,w}$, which is locally $\varepsilon$-LDP on $\mathcal{H}^{\mathrm{hyp}}$, also preserves $\varepsilon$-LDP globally when extended to the joint space $\mathcal{H}^{\mathrm{test;hyp}}$. Crucially, (19) does not imply (20). Indeed, by [23, Theorem 4], the identity channel on $\mathcal{H}^{\mathrm{test}}$ fails to satisfy differential privacy for any $\varepsilon\geq 0$, so the composition $\mathbb{I}^{\mathrm{test}}\otimes\Lambda^{\mathcal{A}}_{s,w}$ cannot satisfy $\varepsilon$-LDP solely on the basis of (19). Therefore, there is a gap in the argument of [6, Appendix C.7]: the claimed bound (18) does not follow from their stated assumption (19). That is, one must assume (20) instead of (19). Moreover, the upper bound obtained in (18) involves the term $I(\text{test};\text{train})_{\rho^{\mathcal{A}}{(s,w)}}$. To render this bound meaningful, $I(\text{test};\text{train})_{\rho^{\mathcal{A}}{(s,w)}}$ should also be controlled by some function of the security parameter $\varepsilon$, although our evaluations—such as Theorem 2—do satisfy this requirement.
C-B Security condition in [6, Appendix C.7]
Their second part essentially changes their model into the Untrusted Data Processor scenario studied in Section V because on the page $59$ of [6] the authors mention the following:
“Next, we turn our attention to the classical MI term in our generalization bounds. Here, we assume that the learner $\mathcal{A}$ uses an overall $\varepsilon$-LDP POVM. As the POVM $\left\{|s\rangle\langle s|\otimes E_{s}^{\mathcal{A}}(w)\right\}_{s,w}$ is not LDP even if every $\left\{E_{s}^{\mathcal{A}}(w)\right\}_{w}$ is, we make the simplifying assumption that the learner uses an $s$-independent $\varepsilon$-LDP POVM $\left\{E^{\mathcal{A}}(w)\right\}_{w}$.”
Even in this scenario, our results still hold, as explained in Section V. However, in this scenario, it is reasonable to impose the ITA condition given in Definition 7 to our learning algorithm, as discussed in Section V while they did not consider such a constraint.
Appendix D Proof of Theorem 1
Given the fact that $I[S\bm{\mathfrak{Te}};WB^{\prime}]=D(\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})$, we can lower-bound $D(\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})$ as follows,
$~~~D(\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})$
$\overset{(a)}{\geq}\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\log\mathrm{Tr}\left[e^{\lambda L^{S\bm{\mathfrak{Te}}WB}}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]$
$=\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\log\mathrm{Tr}\left[e^{\lambda\left(L^{S\bm{\mathfrak{Te}}WB}-\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]\mathbb{I}^{S\bm{\mathfrak{Te}}WB}\right)}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]$
$~~~-\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]$
$\overset{(b)}{\geq}\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]-\frac{\lambda^{2}\alpha^{2}}{,}{2}$ (21)
where $(a)$ follows from the variational lower-bound for the quantum relative entropy (see [13, Theorem $5.9$]) and $(b)$ follows from (7) of Definition 5. Further, we can rewrite (21) as follows,
$\frac{\lambda^{2}\alpha^{2}}{2}-\lambda\left(\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]\right)$
$\hskip 250.0pt+D(\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\geq 0,$
Since the above inequality is a non-negative quadratic equation in $\lambda$ with the coefficient $\frac{\alpha^{2}}{2}\geq 0,$ its discriminant must be non-positive. Hence,
$\left(\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]\right)^{2}\hskip 100.0pt$
$\leq 4\cdot\frac{\alpha^{2}}{2}\cdot D(\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}).$ (22)
Thus, we have,
$\left|\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]\right|\leq\sqrt{2\alpha^{2}I[S\bm{\mathfrak{Te}};WB^{\prime}]}.$ (23)
Therefore, the combination of Definition 4 and (23) yields the following,
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ $\leq\sqrt{2\alpha^{2}I[S\bm{\mathfrak{Te}};WB^{\prime}]}.$ (24)
This completes the proof of Theorem 1.
Appendix E Proof of Corollary 2
Since $s\overset{k}{\sim}s^{\prime}$, there exists a $k+1$-length sequence $\{s_{i}\}_{i=0}^{k}\subseteq\mathcal{S}$ such that $s_{0}=s$, $s_{k}=s^{\prime}$ and for each $i\in[k]$, $s_{i-1}\overset{1}{\sim}s_{i}$. Thus, for any $0\preceq\Lambda\preceq\mathbb{I}$, using Eq (9), we have
$\mathrm{Tr}[\Lambda\mathcal{N}^{(s)}(\rho_{s})]$ $\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s_{1})}(\rho_{s_{1}})]+\delta$
$\leq e^{2\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s_{2})}(\rho_{s_{2}})]+(e^{\varepsilon}+1)\delta$
$\leq e^{3\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s_{3})}(\rho_{s_{3}})]+(e^{2\varepsilon}+e^{\varepsilon}+1)\delta$
[…]
$\leq e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})]+(e^{(k-1)\varepsilon}+e^{(k-2)\varepsilon}+\cdots+e^{\varepsilon}+1)\delta$
$=e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}(\sigma)]+g_{k}(\delta),$
This completes the proof of Corollary 2.
Appendix F Proofs of Theorem 2, Corollary 3 and Corollary 4
F-A Proof of Theorem 2
The proof of Theorem 2 above relies on the Claims 1 and 2 below. The proofs of these claims are given in Appendix I and J.
Claim 1.
Consider $\rho,\rho^{\prime},\sigma\in\mathcal{D}(\mathcal{H}_{A})$. Then, $D(\rho\|\sigma)\leq D(\rho\|\rho^{\prime})+D_{\max}(\rho^{\prime}\|\sigma).$
Claim 2.
Consider $\rho$ and $\sigma$ be two quantum states over Hilbert space $\mathcal{H}$ such that $\rho\ll\sigma$ and $\sigma$ is a finite mixture of probability distributions such that $\sigma=\sum_{b=1}^{m}P(b)\sigma_{b}$, where $\sum_{b=1}^{m}P(b)=1$, and $\rho\ll\sigma_{b}$ for all $b\in[m]$. Then, $D(\rho\|\sigma)\leq\min_{b\in[m]}\left\{D\left(\rho\|\sigma_{b}\right)-\log P(b)\right\}.$
Since we aim to obtain an upper-bound on $I[S;WB^{\prime}]$, one way to proceed is to use the fact that $I[S;WB^{\prime}]=\min_{\omega^{B}}D(\sigma^{SB}\|\sigma^{S}\otimes\omega^{B})$ (where $B\equiv WB^{\prime}$). Thus,
$I[S;WB^{\prime}]$ $\leq D(\sigma^{SB}\|\sigma^{S}\otimes\omega^{B})$
$=\sum_{s\in\mathcal{S}}P^{\otimes n}_{Z}(s)D(\mathcal{N}^{s}(\rho_{s})\|\omega^{B}).$ (25)
We now choose different values of $\omega^{B}$ to obtain upper-bounds on $I[S;WB^{\prime}]$ discussed in steps below.
(Step $\bf 1$}) Consider $\omega^{B}$ to be a uniform mixture of $\mathcal{N}^{\bf f}(\rho_{\mathbf{f}}),$ over all the types $\mathbf{f}\in T^{n}_{\left|\mathcal{Z}\right|}$ i.e. $\omega^{B}:=\frac{1}{\left|T^{n}_{\left|\mathcal{Z}\right|}\right|}\sum_{\mathbf{f}\in T^{n}_{\left|\mathcal{Z}\right|}}\mathcal{N}^{\bf f}(\rho_{\mathbf{f}})$, then, using Claim 2 and the fact that $\left|T^{n}_{\left|\mathcal{Z}\right|}\right|\leq(n+1)^{\left|\mathcal{Z}\right|-1}$ (see [12, Eq. $6.18$]), Eq. (25) can be upper-bounded as follows,
$~~~I[S;WB^{\prime}]$
$\leq\sum_{s\in\mathcal{S}}P^{\otimes n}_{Z}(s)\min_{\mathbf{f}\in T^{n}_{\left|\mathcal{Z}\right|}}\left\{D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{\bf f}(\rho_{\mathbf{f}}))-\log\left|T^{n}_{\left|\mathcal{Z}\right|}\right|^{-1}\right\}$
$\leq(\left|\mathcal{Z}\right|-1)\log(n+1).$
Observe that the above upper-bound on $I[S;WB^{\prime}]$ is independent of the privacy parameters ($\varepsilon,\delta$) of $\mathcal{A}$. This happened because, we chose $\omega^{B}$ to be a uniform mixture of representative quantum states of each type. This choice implied that $\min_{\mathbf{f}\in T^{n}_{\left|\mathcal{Z}\right|}}D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{\bf f}(\rho_{\mathbf{f}}))=0$. To get an upper-bound on $I[S;WB^{\prime}]$ in terms of the privacy parameters, we need to choose $\omega^{B}$ which makes use of the fact that $\mathcal{A}$ satisfies (9). We will accomplish this by using a grid covering for the types of $\mathcal{S}$. We discuss in the step below.
(Step $\bf 2$}) In contrast to Step $1$, we will now choose $\omega^{B}$ to be a mixture over a smaller collection of the output states of $\mathcal{A}$. This smaller collection is obtained by using a grid covering over the types of $\mathcal{S}$, which was developed in the proof of Proposition $2$ of [17]. We now discuss their grid covering over the types of $\mathcal{S}$ below.
Observe that any type ${\bf f}\in T^{n}_{\left|\mathcal{Z}\right|}$ can be thought of as a point inside a $\left|\mathcal{Z}\right|-1$ dimensional grid $0,n]^{\left|\mathcal{Z}\right|-1}$, which is of size $(n+1)^{\left|\mathcal{Z}\right|-1}$. This is because, for any ${\bf f}=({\bf f}_{1},\cdots,{\bf f}_{\left|\mathcal{Z}\right|})\in T^{n}_{\left|\mathcal{Z}\right|}$, the first $\left|\mathcal{Z}\right|-1$ coordinates decide the last coordinate ${\bf f}_{\left|\mathcal{Z}\right|}$, since we have a constraint $\sum_{i=1}^{\left|\mathcal{Z}\right|}{\bf f}_{i}=n$. We now split each dimension of the grid $0,n]^{\left|\mathcal{Z}\right|-1}$ (which is a $[0,n]$ interval) into $t$ equal parts for some
$t\in\mathbb{N}:t\in[1,n].$ (26)
We can think of the grid $0,n]^{\left|\mathcal{Z}\right|-1}$ as a cover of $t^{\left|\mathcal{Z}\right|-1}$ smaller grids of length $l:=\frac{n}{t}$. Note that each side of the smaller grid has $\left\lfloor l\right\rfloor+1$ points. Further, if $\left\lfloor l\right\rfloor+1$ is odd, then we choose the central point of the smaller corresponding to the coordinates of the center of the smaller grid. Thus, for any $s\in\mathcal{S}$, if we consider its type as ${\bf f}^{(s)}$, then we can find a type ${\bf g}^{(s)}\in T^{n}_{\left|\mathcal{Z}\right|}$ such that the first $\left|\mathcal{Z}\right|-1$ coordinates of $\bf f$} are the coordinates of the center of the smaller grid in which the first $\left|\mathcal{Z}\right|-1$ coordinates of ${\bf f}^{(s)}$ resides. In each dimension of the bigger grid, the distance between $s$ and the center of the nearest smaller grid ${\bf c}^{s}$ is given as follows,
$\left|{\bf f}^{(s)}_{z_{(i)}}-{\bf c}^{(s)}_{z_{(i)}}\right|\leq\frac{\left\lfloor l\right\rfloor+1}{2}\leq\frac{n}{2t}+\frac{1}{2},\mbox{~for each }i\in[\left|\mathcal{Z}\right|-1],$
where $z_{(i)}$ is the $i$-th element of the alphabet $\mathcal{Z}$. Therefore, if along all dimension $i\in[\left|\mathcal{Z}\right|-1]$, ${\bf f}^{(s)}_{z_{(i)}}-{\bf g}^{(s)}_{z_{(i)}}=-\frac{n}{2t}+\frac{1}{2}$, then the count of last element $z_{(\left|\mathcal{Z}\right|)}\in\mathcal{Z}$ has to compensate for it. Thus, we have the following,
$\left|{\bf g}^{(s)}_{z_{(\left|\mathcal{Z}\right|)}}-{\bf f}_{z_{(\left|\mathcal{Z}\right|)}}\right|$ $\leq(\left|\mathcal{Z}\right|-1)\left(\frac{n}{2t}+\frac{1}{2}\right).$
Then, for any $s\in\mathcal{S}$ the following holds,
$\min_{s^{\prime}\in T_{{\bf g}^{(s)}}}d(s,s^{\prime})$ $\leq(\left|\mathcal{Z}\right|-1)\frac{n}{t},$ (27)
where $d(s,s^{\prime})$ is the Hamming distance between $s$ and $s^{\prime}$.
(Step $\bf 3$}) We now prove Theorem 2 using the grid covering technique discussed in the proof of [17, Proposition 2]. Fix $\omega^{B}=\sum_{{\bf f}\in T^{\prime}}\frac{1}{\left|T^{\prime}\right|}\mathcal{N}^{\bf f}(\rho_{\bf f})$, where $T^{\prime}$ is the collection of the center points of all the smaller grids obtained in Step 2. Then, using Claim 2 and the fact that $\left|T^{\prime}\right|\leq t^{\left|\mathcal{Z}\right|-1}$, we have,
$I[S;WB^{\prime}]$ $\leq\sum_{s\in\mathcal{S}}P^{\otimes n}_{Z}(s)\min_{{\bf f}\in T^{\prime}}\left\{D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{\bf f}(\rho_{\bf f}))+(\left|\mathcal{Z}\right|-1)\log t\right\}$
$\leq\sum_{s\in\mathcal{S}}P^{\otimes n}_{Z}(s)\left(D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{{\bf g}^{(s)}}(\rho_{{\bf g}^{(s)}}))+(\left|\mathcal{Z}\right|-1)\log t\right).$ (28)
(Step $\bf 4$}) We will now analyze the first term in the RHS of (28) by using Claim 1 and [18, Lemma $6.9$]. Toward this, in [18, Lemma $6.9$], let $\rho=\mathcal{N}^{s}(\rho_{s})$ and $\sigma=\mathcal{N}^{{\bf g}^{(s)}}(\rho_{{\bf g}^{(s)}})$. Thus, [18, Lemma $6.9$] implies that there exists a quantum state $\mathcal{N}^{s}(\rho_{s})^{\prime}$ in the close vicinity of $\mathcal{N}^{s}(\rho_{s})$ such that $D_{\max}(\mathcal{N}^{s}(\rho_{s})^{\prime}\|\mathcal{N}^{{\bf g}^{(s)}}(\rho_{{\bf g}^{(s)}}))\leq f(\varepsilon,\delta)$, where $f(\cdot,\cdot)$ is some function. Thus, using [18, Lemma $6.9$], Claim 1, Assumption (11) and the extension of privacy constraints of $\mathcal{A}$ under $k$-neighboring inputs, we have the following series of inequalities,
$D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{{\bf g}^{(s)}}(\rho_{{\bf g}^{(s)}}))$ $\leq D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{s}(\rho_{s})^{\prime})+\varepsilon^{\prime}$
$\leq\frac{2}{m}E^{2}_{1}\left(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{s}(\rho_{s})^{\prime}\right)+\varepsilon^{\prime}$
$\leq\varepsilon^{\prime}+\frac{2}{m}g_{\frac{n(\left|\mathcal{Z}\right|-1)}{t}}(\varepsilon,\delta),$ (29)
where $\varepsilon^{\prime}:=\frac{n(\left|\mathcal{Z}\right|-1)\varepsilon}{t}+\log\frac{1}{1-g_{\frac{n(\left|\mathcal{Z}\right|-1)}{t}}(\varepsilon,\delta)}$, and $g_{\frac{n(\left|\mathcal{Z}\right|-1)}{t}}(\varepsilon,\delta)=\frac{e^{\frac{n(\left|\mathcal{Z}\right|-1)\varepsilon}{t}}-1}{e^{\varepsilon}-1}\delta$. Thus, using Eqs. (28) and (29) we have,
$\hskip-7.0ptI[S;WB^{\prime}]{\leq}\frac{n(\left|\mathcal{Z}\right|-1)\varepsilon}{t}+(\left|\mathcal{Z}\right|-1)\log t+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta),$ (30)
where $h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta):=\log\frac{1}{1-g_{{n(\left|\mathcal{Z}\right|-1)}}(\delta)}+\frac{2}{m}g_{n(\left|\mathcal{Z}\right|-1)}(\delta)$ (observe that $h_{\left|\mathcal{Z}\right|}(\varepsilon,0)=0$) and the last inequality follows from the fact that the grid size $t\geq 1$.
(Step $\bf 5$}) In this step, we optimize the choice over $t$ (grid size) to tighten the upper-bound obtained in Eq. (30). Observe that the value of $t$ which minimizes the RHS of Eq. (30) is,
$t^{\star}=n\varepsilon.$ (31)
As mentioned in the statement of Theorem 2, we have $\varepsilon\in\left[\frac{1}{n}\right.,1]$ and thus (31) yields that $1\leq t^{\star}\leq n$, which satisfies the size constraint of grid mentioned in (26). Therefore, substituting $t=t^{\star}$ in (30) yields,
$I[S;WB^{\prime}]$ $\leq(|\mathcal{Z}|-1)\left(1+\log\left({n\varepsilon}\right)\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)$
$=(|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$
This completes the proof of Theorem 2.
F-B Proof of Corollary 3
For $\varepsilon<\frac{1}{n}$ (as mentioned in Corollary 3), (31) yields that $t^{\star}<1$, and therefore it does not (26). Thus, in this case, we set $t=1$ in (30) to obtain the desired upper-bound.
F-C Proof of Corollary 4
For $\varepsilon>1$ (as mentioned in Corollary 4), (31) yields that $t^{\star}>n$, which does not satisfy the grid size constraint mentioned in (26). Therefore in this case we set grid size $t=n$. However, for this choice of grid size, observe that the grid covers all the sequences in $\mathcal{S}$ and therefore covers all the type-representatives in $\mathcal{S}$. This is the same case as Step 1. Therefore, we have,
$I[S;WB^{\prime}]$ $\leq(\left|\mathcal{Z}\right|-1)\log(n+1).$
This completes the proof of Corollary 4. Further, note that if we substitute $t=n$ in (30), then it would yield us a weaker bound as compared to the above.
Appendix G Proof of Lemma 1
Choose a basis $\{|x\rangle\}$ that diagonalizes $\rho_{s}$. In this basis, the state can be expressed as
$\rho_{s}=\sum_{x}P_{X|s}(x)\,|x\rangle\langle x|,$ (32)
on system $X$. Next, define the instrument $\{\mathcal{N}_{w}^{\prime}\}_{w}$ by
$\mathcal{N}_{w}^{\prime}(|x\rangle\langle x|):=|x\rangle\langle x|\otimes\mathcal{N}_{w}(|x\rangle\langle x|),$ (33)
where the output system is $XB^{\prime}$. Recall that the original learning algorithm $\{\mathcal{N}_{w}\}_{w}$ outputs on system $B^{\prime}$. Under this construction, $\{\mathcal{N}_{w}^{\prime}\}_{w}$ is strictly more informative than $\{\mathcal{N}_{w}\}_{w}$.
Suppose, for contradiction, that $\{\mathcal{N}_{w}\}_{w}$ is more informative than $\{\mathcal{N}_{w}^{\prime}\}_{w}$. Then there exist CP-TP maps $\{\Gamma_{w}\}_{w}$ such that
$\Gamma_{w}(\mathcal{N}_{w}(\rho_{s}))=\mathcal{N}_{w}^{\prime}(\rho_{s})\quad\text{for all }s\in\mathcal{S}.$
Hence,
$\mathrm{Tr}_{B^{\prime}W}\Bigg[\sum_{w}\Gamma_{w}(\mathcal{N}_{w}(\rho_{s}))\otimes|w\rangle\langle w|\Bigg]$
$=$ $\mathrm{Tr}_{B^{\prime}W}\Bigg[\sum_{w}\mathcal{N}_{w}^{\prime}\Bigg(\sum_{x}P_{X|s}(x)|x\rangle\langle x|\Bigg)\otimes|w\rangle\langle w|\Bigg]$
$=$ $\mathrm{Tr}_{B^{\prime}W}\Bigg[\sum_{x}P_{X|s}(x)\sum_{w}\mathcal{N}_{w}^{\prime}(|x\rangle\langle x|)\otimes|w\rangle\langle w|\Bigg]$
$=$ $\mathrm{Tr}_{B^{\prime}W}\Bigg[\sum_{x}P_{X|s}(x)|x\rangle\langle x|\otimes\sum_{w}\mathcal{N}_{w}(|x\rangle\langle x|)\otimes|w\rangle\langle w|\Bigg]$
$=$ $\sum_{x}P_{X|s}(x)|x\rangle\langle x|=\rho_{s},$ (34)
which contradicts the assumption that no CP-TP map $\Gamma$ satisfies
$\Gamma\left(\sum_{w\in\mathcal{W}}\mathcal{N}_{w}(\rho_{s})\otimes|w\rangle\langle w|\right)=\rho_{s}.$
Therefore, $\{\mathcal{N}_{w}^{\prime}\}_{w}$ is strictly more informative than $\{\mathcal{N}_{w}\}_{w}$, and thus $\{\mathcal{N}_{w}\}_{w}$ is not ITA.
Figure 3: Numerical comparison of the generalization error bounds for the classical-quantum toy example in [19]. (a) Comparison of $\mathcal{B}_{\text{MI}}$ (35) and $\mathcal{B}_{\text{SEP}}$ (36) as a function of the prior probability $p\in[0.25,0.75]$. (b) Comparison of $\mathcal{B}_{\text{MI}}$ (35) and $\mathcal{B}_{\text{SEP}}$ (36) as a function of the sub-Gaussianity parameter $\alpha\in[0.1,1]$ for a fixed prior $p=0.4$. In both regimes, our bound $\mathcal{B}_{\text{MI}}$ (blue) provides a strictly tighter upper bound than $\mathcal{B}_{\text{SEP}}$ (orange).
Appendix H Numerical Comparison of Generalization Error Bounds
In this appendix, we numerically validate our theoretical results by comparing them against the bounds established in [19]. We utilize the classical-quantum toy example described in [19, Section VI] to demonstrate the tightness of our mutual information-based approach.
For this comparison, we evaluate the following two quantities under the condition that the sub-Gaussianity parameters satisfy $\mu=\tau=\alpha$:
  1. 1.
    Our Mutual Information Bound ($\mathcal{B}_{\text{MI}}$): Derived from Theorem 1, this bound relies on the total mutual information between the input and the output system. Due to the independence of the test and train systems conditioned on $Z$ in this example, the term $I(Z,\bm{\mathfrak{Te}};WB^{\prime})$ simplifies, yielding:
    ${\color[rgb]{0.12109375,0.46484375,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.46484375,0.70703125}\mathcal{B}_{\text{MI}}=\sqrt{2\alpha^{2}I(Z;WB^{\prime})}}.$ (35)
  2. 2.
    The Separated Bound from [19] ($\mathcal{B}_{\text{SEP}}$): We compare against the bound in [19, Eq. (88), Theorem 1], which separates the classical and quantum contributions. In this specific toy example, the first term of their theorem vanishes, reducing the bound to:
    ${\color[rgb]{1,0.49609375,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49609375,0.0546875}\mathcal{B}_{\text{SEP}}=\mathbb{E}_{Z,W}\left[\sqrt{2\alpha^{2}D\left(\sigma^{B^{\prime}}_{(Z,W)}\middle\|\sigma^{B^{\prime}}_{W}\right)}\right]+\sqrt{2\alpha^{2}I(Z;W)}}.$ (36)
As detailed in Figure 3 above, for comparison, we plot (35) and (36) for the example mentioned in [19, Subsection-VI].
Appendix I Proof of Claim 1
By the definition of the max-relative entropy, we have
$\rho^{\prime}\leq e^{D_{\max}(\rho^{\prime}\|\sigma)}\,\sigma.$
Equivalently,
$\sigma\geq e^{-D_{\max}(\rho^{\prime}\|\sigma)}\,\rho^{\prime}.$
Since the logarithm is operator monotone, this implies
$\log\sigma\succeq\log\rho^{\prime}-D_{\max}(\rho^{\prime}\|\sigma)\,\mathbb{I}.$
Multiplying both sides by $\rho$ and taking the trace (which reverses the inequality), we obtain
$\mathrm{Tr}[\rho\log\sigma]\leq-\mathrm{Tr}[\rho\log\rho^{\prime}]+D_{\max}(\rho^{\prime}\|\sigma).$
Adding $\mathrm{Tr}[\rho\log\rho]$ to both sides gives
$\mathrm{Tr}[\rho(\log\rho-\log\sigma)]\leq\mathrm{Tr}[\rho(\log\rho-\log\rho^{\prime})]+D_{\max}(\rho^{\prime}\|\sigma),$
which can be written as
$D(\rho\|\sigma)\leq D(\rho\|\rho^{\prime})+D_{\max}(\rho^{\prime}\|\sigma).$
This completes the proof of Claim 1.
Appendix J Proof of Claim 2
We begin by invoking the operator monotonicity of the function $\log(\cdot)$. Since $\sigma\succeq P(b)\sigma_{b}$ for every $b$, we obtain
$\log\sigma\succeq\log P(b)\,\mathbb{I}+\log\sigma_{b}.$ (37)
Using (37), we immediately have,
$D(\rho\|\sigma)$ $=\mathrm{Tr}\left[\rho(\log\rho-\log\sigma)\right]$
$\leq\mathrm{Tr}\left[\rho(\log\rho-\log P(b)\,\mathbb{I}-\log\sigma_{b})\right]$
$=D(\rho\|\sigma_{b})-\log P(b).$ (38)
Since (38) holds for all $b$, it implies Claim 2.