### v2

# Equivalence of Privacy and Stability with Generalization Guarantees in Quantum Learning

Abstract
We present a unified information-theoretic framework elucidating the interplay between stability, privacy, and the generalization performance of quantum learning algorithms. We establish a bound on the expected generalization error in terms of quantum mutual information and derive a probabilistic upper bound that generalizes the classical result by Esposito et al. (2021). Complementing these findings, we provide a lower bound on the expected true loss relative to the expected empirical loss.
Additionally, we demonstrate that $(\varepsilon,\delta)$-quantum differentially private learning algorithms are stable, thereby ensuring strong generalization guarantees. Finally, we extend our analysis to dishonest learning algorithms, introducing Information-Theoretic Admissibility (ITA) to characterize the fundamental limits of privacy when the learning algorithm is oblivious to specific dataset instances.
I Introduction
I-A Background and Motivation (Classical to Quantum)
A central goal of statistical learning theory is to understand when a hypothesis trained on finite data generalizes to unseen examples. In the classical domain, a profound equivalence exists between algorithmic stability—the insensitivity of a model to training set perturbations—and its generalization performance [5]. Concurrently, Differential Privacy (DP) [12] has emerged as the rigorous standard for stability in randomized algorithms, preventing overfitting even in adaptive data analysis [10, 2, 23].
As Quantum Machine Learning (QML) matures [4], it becomes urgent to establish classical-style guarantees—such as stability-to-generalization implications and privacy-based stability—for learning from quantum data [8, 26]. While Quantum DP (QDP) [30] successfully extends indistinguishability to quantum states, the intersection of QDP with statistical learning theory remains under-explored. The non-commutative nature of quantum information and the irreversibility of measurement necessitate a rigorous information-theoretic treatment.
I-B Our Framework and Main Contributions
In this work, we present a unified framework to determine the fundamental limits of generalization in quantum learning algorithms subject to privacy constraints. We adopt an information-theoretic approach, utilizing the mutual information between the training data and the algorithm’s output as the primary metric for stability [29].
To operationalize this study, we frame the learning protocol as an interaction between three distinct parties. The Respondent is the collection of individuals who contribute their private data to the training set; their primary concern is to prevent the leakage of their specific data instances. The Data Processor is the entity that aggregates this data and executes the quantum learning algorithm. Finally, the Investigator is the end-user who receives the algorithm’s output (the hypothesis) and whose goal is to minimize generalization error on unseen data. In the standard trusted setting, the Data Processor acts to protect data from the Investigator; in the untrusted setting, the Respondents require protection even from the Data Processor.
We structure our analysis around three pivotal logical steps: first establishing that stability governs generalization, then demonstrating that privacy enforces this stability, and finally extending these concepts to untrusted environments via Information-Theoretic Admissibility.
First step: Stability implies Generalizability. We first address the Investigator’s goal of generalization. Building on the classical equivalence between uniform stability and generalization [5], we prove that stable quantum algorithms inherently generalize. Specifically, by assuming a classical-quantum sub-Gaussian property for the loss function, we derive a bound on the expected generalization error in terms of the mutual information between the training data and the output. This confirms that limiting the algorithm’s dependence on individual quantum data points effectively prevents overfitting. Going further, we prove a bound on generalization error in probability using the Sandwiched Rényi divergence [19] to ensure robust performance guarantees. This bound is the quantum version of [13, Corollary $2$]. Finally, we complement these upper bounds with a lower bound on the expected true loss in terms of the empirical loss, establishing a tight bidirectional relationship between observed and actual risk.
Second step: Privacy implies Generalizability via Stability. Next, we address the Respondent’s privacy via Differential Privacy, which acts as a rigorous form of stability. Extending prior results in classical [11] and pure quantum DP [22], our main contribution is to generalize this link to the $(\varepsilon,\delta)$-DP setting. We propose a "1-neighbor" $(\varepsilon,\delta)$-DP framework and derive a rigorous upper bound on the mutual information between the training data and the quantum output. By combining this with our first step, we formally demonstrate that privacy is a sufficient condition for generalization in quantum learning.
Third step: Dishonest Third Party and Information-Theoretic Admissibility. We extend our framework to the untrusted Data Processor scenario, where the learning map must remain independent of specific training instances to prevent leakage to the processor. To rigorously characterize the limits of this setting, we introduce Information-Theoretic Admissibility (ITA). An algorithm is ITA if it is information-theoretically optimal; specifically, no other algorithm exists that is strictly more informative (allowing superior state recovery via a CP-TP map). Non-ITA algorithms render privacy definitions ineffective, as an adversary could employ a superior extraction strategy. Crucially, we demonstrate a fundamental separation: while classical ITA implies full data recoverability (precluding privacy), the quantum setting permits non-trivial ITA algorithms where privacy remains meaningful. This highlights a genuine quantum advantage, validating our security definitions even under the constraint of optimal information extraction.
I-C Our Contributions
Our work comprehensively characterizes the interplay between privacy, stability, and generalization in quantum learning:
  • Expected Generalization: We introduce Classical-Quantum Sub-Gaussianity to bound the expected generalization error by the square root of mutual information (Theorem 1). This unifies classical and quantum fluctuations, confirming stability as an equivalent condition for generalizing capability.
  • Probabilistic Bound: We establish an upper-bound on the generalization error in probability (Theorem 2) using the Sandwiched Rényi Divergence assuming an i.i.d. structure of the data and algorithm. Our unified framework leverages the non-commutative Hölder inequality to demonstrate an $O(1/\sqrt{n})$ convergence rate of the generalization error.
  • Lower Bound on Expected True Loss: We derive a reverse inequality linking expected true loss to empirical loss (Theorem 3). This establishes a tight relationship between empirical observations and true risk, recovering classical analogues [18] as limiting cases.
  • Stability of $(\varepsilon,\delta)$-QDP Learning Algorithms: We derive a mechanism-agnostic upper bound on the mutual information for $1$-neighbor $(\varepsilon,\delta)$-DP quantum learning algorithms (Theorem 4). Using grid-covering techniques, we generalize classical results [23] to show that privacy enforces stability with logarithmic sample size scaling.
  • Untrusted Processor Framework: We introduce Information-Theoretic Admissibility (ITA) for untrusted settings. We prove a fundamental quantum advantage: unlike classical settings where ITA precludes privacy, quantum non-commutativity allows for optimal algorithms that simultaneously maintain differential privacy.
II Notations
Let $\mathcal{D}(\mathcal{H})$ denote the set of density operators on a finite-dimensional Hilbert space $\mathcal{H}$.
Method of Types.
[…]
III A General Framework for Quantum Learning Algorithms
III-A Learning Setup and Data Encoding
In this section, we establish a quantum learning framework motivated by [8] and [26], operationalized through the interaction between a Respondent (data contributor) and a Data Processor (algorithm executor), focusing on the non-private setting (i.e., without imposing any privacy constraint at this stage).
[…]
The Data Processor receives the classical-quantum input $\sum_{s}P_{S}(s)|s\rangle\langle s|\otimes\rho_{s}$ and executes a learning algorithm, modeled as a collection of quantum instruments $\mathcal{N}:=\{\mathcal{N}^{(s)}:\mathcal{H}^{\bm{\mathfrak{Tr}}}\to\mathcal{H}^{B}\}_{s\in\mathcal{S}}$. The output system $B\equiv WB^{\prime}$ comprises a classical hypothesis $W$ and a quantum residue $B^{\prime}$. The resulting joint state is given by:
$\sigma^{S\bm{\mathfrak{Te}}B}_{\mathcal{N}}:=\sum_{s\in\mathcal{S}}P_{S}(s)|s\rangle\langle s|\otimes(\sigma^{\mathcal{N}}_{s})^{\textnormal{Te}B},$ (1)
where the output state conditioned on the Respondents’ input $s$ is,
$\sigma^{\mathcal{N}}_{s}:=\sum_{w\in\mathcal{W}}((\mathbb{I}^{\bm{\mathfrak{Te}}}\otimes\mathcal{N}^{(s)}_{w})(\rho_{s}))^{\bm{\mathfrak{Te}}B^{\prime}}\otimes|w\rangle\langle w|^{W}.$ (2)
Here, the Data Processor’s output system $B$ comprises a classical hypothesis system $W$ and a quantum residual system $B^{\prime}$, i.e., $B\equiv WB^{\prime}$. The action of the Data Processor’s instrument $\mathcal{N}^{(s)}$ is denoted as:
$\mathcal{N}^{(s)}(\rho_{s}):=\sum_{w\in\mathcal{W}}(\mathbb{I}^{\bm{\mathfrak{Te}}}\otimes\mathcal{N}^{(s)}_{w})(\rho_{s})\otimes|w\rangle\langle w|,$ (3)
where each $\mathcal{N}^{(s)}_{w}$ is a completely positive trace non-increasing map. This interaction is illustrated in Figure 1.
$\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$Respondent $\mathbf{\sum_{s}P_{S}(s)|s\rangle\langle s|\otimes\rho_{s}}$ $\mathbf{\{s,\rho_{s}\}}$$\mathbf{\mathcal{N}^{(s)}(\rho_{s})}$Learning AlgorithmData Processor
Figure 1: Privacy based learning framework.
III-B Stability of a Quantum Learning Algorithm
We now define stability for quantum learning algorithms. Intuitively, stability requires the learning outcome to remain invariant to single-entry modifications, thereby preventing the leakage of individual data points. Extending the classical information-theoretic framework of [29], which quantifies stability via mutual information, we formalize this notion below.
Definition 1.
(Stability) A quantum learning algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s}$ is defined to be $\gamma$-stable, if $\max_{P_{S}}I\left[S\bm{\mathfrak{Te}};WB^{\prime}\right]\leq\gamma$, where $I\left[S\bm{\mathfrak{Te}};WB^{\prime}\right]$ is calculated with respect to the classical-quantum state mentioned in (2).
The above definition provides a quantitative upper bound on the information that can be extracted from the Data Processor’s output $B$ ($B\equiv WB^{\prime}$) about the Respondent’s input dataset $S$ and $\bm{\mathfrak{Te}}$. Consequently, a small upper-bound implies that the algorithm’s output is not strongly dependent on any single training data point, indicating that the algorithm is information-theoretically stable.
III-C Stability Implies Generalizability For Quantum Learning Algorithms
In this section, we demonstrate that if the Data Processor employs a stable algorithm, the results generalize well to unseen data. For a quantum learning algorithm $\mathcal{N}=\{\mathcal{N}^{(s)}\}$, the joint state representing the Respondent’s input and the Data Processor’s output, mentioned in (1), can be expanded as:
$\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}:=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}P_{S}(s)\overset{S}{|s\rangle\langle s|}\otimes P^{\mathcal{N}}_{W\mid S}(w\mid s)\overset{W}{|w\rangle\langle w|}\otimes(\sigma^{\mathcal{N}}_{s,w})^{\bm{\mathfrak{Te}}B^{\prime}},$ (4)
where $P^{\mathcal{N}}_{W\mid S}(w\mid s):=\mathrm{Tr}\left[(\mathbb{I}^{\bm{\mathfrak{Te}}}\otimes\mathcal{N}^{(s)}_{w})(\rho_{s})\right]$ is the probability of the Data Processor selecting hypothesis $w$ given the dataset $s$, and $\sigma^{\mathcal{N}}_{s,w}$ is the normalized residual state $\frac{(\mathbb{I}^{\bm{\mathfrak{Te}}}\otimes\mathcal{N}^{(s)}_{w})(\rho_{s})}{P^{\mathcal{N}}_{W\mid S}(w\mid s)}$. In the following discussion, we define how to quantize the loss or error induced from the resultant state $\sigma^{\mathcal{N}}$.
In the earlier discussed quantum learning framework, the input data $s$ and output hypothesis $w$ induced by the quantum learning algorithm $\mathcal{N}$ are embedded into the output residue quantum state $\sigma^{\mathcal{N}}_{s,w}$. Therefore, to evaluate the performance of the Data Processor, we define the loss in terms of the expected value of observables with respect to the state $\sigma^{\mathcal{N}}$ produced by the Data Processor. In [8, 26], the authors consider a family of non-negative self-adjoint loss observables $\left\{L(s,w)\right\}_{\begin{subarray}{c}(w,s)\end{subarray}}$ which act on the quantum testing system $\bm{\mathfrak{Te}}$ and the output quantum system $B^{\prime}$. Using these loss observables, we define the following global loss operator,
$L^{S\bm{\mathfrak{Te}}WB^{\prime}}:=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}\overset{S}{|s\rangle\langle s|}\otimes\overset{W}{|w\rangle\langle w|}\otimes\overset{\bm{\mathfrak{Te}}B^{\prime}}{L(s,w)}.$ (5)
Based on the above description of the joint state $\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}$ and the loss operators $\{L(s,w)\}$, we now distinguish between the loss observed by the Data Processor on the training data (empirical) and the loss expected on unseen fresh data (true).
Definition 2 (Expected Empirical Loss [8, Definition $11$]).
The expected empirical loss $\hat{L}_{\rho}(\mathcal{N})$ captures the performance of the Data Processor’s algorithm on the dataset provided by the Respondent. It is the expectation over the joint distribution induced by the algorithm:
$\hat{L}_{\rho}(\mathcal{N})$ $:=\mathbb{E}_{(S,W)\sim{P}^{\mathcal{N}}_{SW}}[\mathrm{Tr}[L(S,W)(\sigma^{\mathcal{N}}_{S,W})^{\bm{\mathfrak{Te}}B^{\prime}}]]=\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB^{\prime}}\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}].$
Definition 3 (Expected True Loss [26, Definition $19$]).
The expected true loss
$L_{\rho}(\mathcal{N})$ represents the generalization performance of the Data Processor’s algorithm. It evaluates the hypothesis $W$ generated by the Data Processor against a fresh dataset $\overline{S}$ independent of the training data $S$:
$L_{\rho}(\mathcal{N})$ $:=\mathbb{E}_{(\overline{S},\overline{W})\sim P_{S}\times{P}^{\mathcal{N}}_{{W}}}\left[\mathrm{Tr}\left[L(\overline{S},\overline{W})\left(\rho^{\bm{\mathfrak{Te}}}_{\overline{S}}\otimes(\sigma^{\mathcal{N}}_{\overline{W}})^{B^{\prime}}\right)\right]\right]=\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB^{\prime}}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})].$
where for any $s$, we define $\rho^{\bm{\mathfrak{Te}}}_{s}:=\mathrm{Tr}_{\bm{\mathfrak{Tr}}}[\rho_{s}]$, for each $w$, we define $\sigma^{\mathcal{N}}_{w}:=\mathbb{E}_{S\sim P^{\mathcal{N}}_{S|W=w}}[\mathrm{Tr}_{\bm{\mathfrak{Te}}}[\sigma^{\mathcal{N}}_{S,w}]]$, and $\sigma^{S\bm{\mathfrak{Te}}}$ and $\sigma^{WB^{\prime}}_{\mathcal{N}}$ are the corresponding marginals of the state $\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}$ defined in (4).
Remark 1.
We adopt the definition of expected true loss as proposed in [26] and not that of [8]. The authors in [26] give a rigorous justification for Definition 3 and argue that the definition proposed by [8, Definition $12$] is not a correct definition for the expected true loss.
Based on these definitions, the expected generalization error is defined as the deviation between the Data Processor’s empirical performance and the true performance.
Definition 4 (Expected Generalization Error [26]).
The expected generalization error is:
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ $:=\left|\hat{L}_{\rho}(\mathcal{N})-L_{\rho}(\mathcal{N}))\right|=\left|\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}]-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\right|.$
We will now bound $\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ in terms of $I\left[S\bm{\mathfrak{Te}};WB^{\prime}\right].$ To obtain such a bound in the classical setting [29] assumed that the loss function is sub-Gaussian. We will make a similar assumption for the loss operators $\{L(w,s)\}$ and the Data Processor’s output state. Towards this we make the following definition.
Definition 5 ( Classical-Quantum $\alpha$-Sub-Gaussianity).
For a fixed parameter $\alpha\in(0,\infty),$ the collection $\{L(w,s)\}$ of loss operators is said to be an $\alpha$-sub-Gaussian collection with respect to $\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}:=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}P_{S}(s)|s\rangle\langle s|\otimes P^{\mathcal{N}}_{W}(w\mid s)|w\rangle\langle w|\otimes\rho^{\bm{\mathfrak{Te}}}_{s}\otimes(\sigma^{\mathcal{N}}_{w})^{B^{\prime}},$ if for every $\lambda\in\mathbb{R},$ it satisfies,
$\mathbb{E}\left[\mathrm{Tr}\left[e^{\lambda\left(L(S,W)-\mathbb{E}\left[\mathrm{Tr}\left[L(S,W)\left(\rho^{\bm{\mathfrak{Te}}}_{S}\otimes(\sigma^{\mathcal{N}}_{W})^{B^{\prime}}\right)\right]\mathbb{I}^{\bm{\mathfrak{Te}}B^{\prime}}\right]\right)}\left(\rho^{\bm{\mathfrak{Te}}}_{S}\otimes(\sigma^{\mathcal{N}}_{W})^{B^{\prime}}\right)\right]\right]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2}},$ (6)
where the expectations are calculated with respect to the product distribution $P_{S}\times{P}^{\mathcal{N}}_{{W}}$. Note that (6) is equivalent to,
$\mathrm{Tr}\left[e^{\lambda\left(L^{S\bm{\mathfrak{Te}}WB}-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\mathbb{I}^{S\bm{\mathfrak{Te}}WB}\right)}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2}},$ (7)
where $L^{S\bm{\mathfrak{Te}}WB^{\prime}}$ is the global loss operator defined in (5).
Definition 5 naturally generalizes classical sub-Gaussianity. In the limit of trivial quantum systems ($\dim(\bm{\mathfrak{Te}})=\dim(B^{\prime})=1$), the operators $\sigma^{\mathcal{N}}_{W}$ and $\mathbb{I}^{B^{\prime}}$ become scalars, reducing $L(S,W)$ to a classical random loss function. Consequently, condition (6) collapses to the standard classical sub-Gaussian inequality $\mathbb{E}_{(S,W)}[e^{\lambda(L(S,W)-\mathbb{E}[L(S,W)])}]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2}}$ with respect to $P_{S}\times{P}^{\mathcal{N}}_{{W}}$. We now present a theorem bounding the expected generalization error for quantum learning algorithms.
Theorem 1.
For a fixed $\alpha\in(0,\infty),$ if the loss operators for a quantum learning algorithm $\mathcal{N}$, satisfy Definition 5, then, we have,
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})\leq\sqrt{2\alpha^{2}I[S\bm{\mathfrak{Te}};WB^{\prime}]}.$ (8)
Proof.
See Appendix A for the proof.
The following corollary of Theorem 1, together with Definition 1, indicates that when the Data Processor employs a stable algorithm, the generalization error remains tightly bounded.
Corollary 1.
If the Data Processor’s learning algorithm $\mathcal{N}:=\{\mathcal{N}^{(s)}\}$ is $\gamma$-stable and the loss operators of $\mathcal{N}$ satisfy Definition 5, for a fixed $\alpha\in(0,\infty),$ then, its generalization error is upper bounded by $\sqrt{2\alpha^{2}\gamma}$.
While bounds on the expected generalization error provide a measure of average performance, robust learning requires guarantees that hold with high confidence for individual realizations of the algorithm. To address this, we prove a quantum version of [13, Corollary $2$] in term of the sandwiched Rényi divergence, derived under the assumption of i.i.d. data and loss observable decompositions.
Theorem 2.
Let $\mathcal{N}$ be a quantum learning algorithm. Assume that the associated collection of loss operators $\{L(s,w)\}$ satisfies the Conditional Classical-Quantum Local Sub-Gaussian condition (Definition 8). For any Sandwiched Rényi divergence order $\gamma>1$ and confidence level $\delta\in(0,1)$, the generalization error is bounded with probability at least $1-\delta$ as:
$\Pr_{(S,W)\sim P^{\mathcal{N}}_{SW}}\left\{\textnormal{gen}_{\rho}(\mathcal{N},S,W)\leq\sqrt{\frac{2\alpha^{2}}{n}\left(\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})+\frac{\gamma}{\gamma-1}\ln\frac{2}{\delta}\right)}\right\}\geq 1-\delta,$
where $\textnormal{gen}_{\rho}(\mathcal{N},S,W)$ is the generalization error random variable (Definition 9) and $\tilde{D}_{\gamma}$ denotes the Sandwiched Rényi divergence (defined in (19)).
Proof.
See Appendix C for the complete derivation, which includes a detailed discussion on the i.i.d. structure of the algorithm and the decomposition of the loss operators.
Complementing these upper bounds in expectation and probablity, we provide the following lower bound on the expected true loss in terms of the expected empirical loss.
Theorem 3.
Let $\mathcal{N}$ be a quantum learning algorithm with loss operators satisfying the Classical-Quantum Sub-Gaussian property (Definition 5) with parameter $\alpha>0$. For any sandwiched Rényi divergence order $\gamma>1$, the expected true loss is lower bounded by the empirical loss in the following exponential form,
$\exp\left(L_{\rho}(\mathcal{N})\right)\geq\hat{L}_{\rho}(\mathcal{N})\exp\left(-\left[\frac{\gamma\alpha^{2}}{2(\gamma-1)}+\frac{\gamma-1}{\gamma}\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\right).$ (9)
Proof.
See Appendix B for the proof.
In Appendix D, we compare our upper-bounds on generalization error (Theorem 1 and Theorem 2) with prior works.
IV Generalization Guarantees for Differentially Private Quantum Learning
IV-A Trusted Setting and One-Neighbor $(\varepsilon,\delta)$-DP
This section examines the connection between privacy and generalization in quantum learning. Building on Section III, which linked information-theoretic stability to generalization, we demonstrate that differential privacy enforces this stability. Using the framework of Figure 2, we introduce the Investigator as the recipient of the output system $B$, generated by a Trusted Data Processor from the Respondent’s raw data ($S,\bm{\mathfrak{Te}},\bm{\mathfrak{Tr}}$).
To prevent the reconstruction of individual entries, the Processor ensures the algorithm satisfies differential privacy, requiring output invariance under single-entry modifications. This constraint is mathematically equivalent to algorithmic stability (Definition 1), confirming privacy as a sufficient condition for generalization. We formalize this indistinguishability requirement below.
[…]
Figure 2: Privacy based learning framework.
Definition 6.
An algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$ is a $1$-neighbor $(\varepsilon,\delta)$-DP support consistent learning algorithm if it satisfies the following conditions:
  1. 1.
    Permutation Invariance: For all $s,s^{\prime}\in\mathcal{S}$ satisfying $T_{s}=T_{s^{\prime}}$, the algorithm satisfies the condition $\mathcal{N}^{(s)}(\rho_{s})=\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})$. This ensures that the algorithm’s output depends solely on the frequency of the data, not its specific ordering. This condition is natural in statistical learning, where the order of training examples is irrelevant to the hypothesis, and it further adds an extra layer of privacy.
  2. 2.
    Privacy: For every $s\overset{1}{\sim}s^{\prime}$ and $0\preceq\Lambda\preceq\mathbb{I}$, the following inequality holds:
    $\begin{split}\mathrm{Tr}[\Lambda\mathcal{N}^{(s)}(\rho_{s})]&\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})]+\delta,\\ \mathrm{Tr}[\Lambda\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})]&\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s)}(\rho_{s})]+\delta.\end{split}$ (10)
  3. 3.
    Support Consistency: For every $s\overset{1}{\sim}s^{\prime}$, the output supports are identical, i.e.,
    $\operatorname{supp}(\mathcal{N}^{(s)}(\rho_{s}))=\operatorname{supp}(\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})).$ (11)
Remark 2.
The support consistency condition (11) is automatically satisfied in the pure differential privacy regime ($\delta=0$).
IV-B Privacy Implies Stability: A Mutual-Information Bound
Definition 6 above implies that privacy guarantees extend to $k$-neighbors, albeit with degraded parameters.
Corollary 2.
If $\mathcal{N}$ satisfies Definition 6, then for any inputs $s\overset{k}{\sim}s^{\prime}$ ($k\geq 1$) and $0\preceq\Lambda\preceq\mathbb{I}$, we have, $\mathrm{Tr}[\Lambda\mathcal{N}^{(s)}(\rho_{s})]\leq e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})]+g_{k}(\varepsilon,\delta),$ where $g_{k}(\varepsilon,\delta):=\frac{e^{k\varepsilon}-1}{e^{\varepsilon}-1}\delta$ is assumed to be strictly less than $1$. The symmetric inequality holds by swapping $s$ and $s^{\prime}$.
Proof.
See Appendix H for the proof.
We now utilize the framework established in Section III to analyze the stability of quantum learning algorithms that satisfy Definition 6. For this analysis, we modify the framework by treating the quantum test data system $\bm{\mathfrak{Te}}$ as trivial (i.e., $\dim(\bm{\mathfrak{Te}})=1$).
Under this modification, the stability measure from Definition 1 simplifies to the mutual information between the training data $S$ and the output system $WB^{\prime}$. Therefore, in the theorem below, we derive an upper bound on $I[S;WB^{\prime}]$ for a quantum $(\varepsilon,\delta)$-differentially private (DP) learning algorithm. This derivation relies on the following assumption regarding the noise parameters $\varepsilon$ and $\delta$,
$g_{n(|\mathcal{Z}|-1)}(\varepsilon,\delta)<1,$ (12)
where for any $k\geq 1$, $g_{k}(\varepsilon,\delta)$ is defined in Corollary 2.
Theorem 4.
For $\varepsilon\in\left[\frac{1}{n},1)\right$., consider $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$ to be a $1$-neighbor $(\varepsilon,\delta)$-DP support consistent learning algorithm (see Definition 6) and satisfies the condition (12). Then, the following holds,
$I[S;WB^{\prime}]\leq(|\mathcal{Z}|-1)\ln\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta),$ (13)
where, $n$ is the length of the training data and for some constant $m\in(0,1],$ $h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta):=\ln\frac{1}{1-g_{{n(\left|\mathcal{Z}\right|-1)}}(\delta)}+\frac{2}{m}g_{n(\left|\mathcal{Z}\right|-1)}(\delta)$ and has a property that $h_{\left|\mathcal{Z}\right|}(\varepsilon,0)=0$.
Proof.
See Appendix E-A for the proof.
[…]
Corollary 3.
For $\varepsilon\in\left.[0,\frac{1}{n}\right)$, consider a learning algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$, which satisfies the properties mentioned in Definition 6 and the condition (12). Then, $I[S;WB^{\prime}]\leq(|\mathcal{Z}|-1)\varepsilon n+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$
Proof.
See Appendix E-B for the proof.
Corollary 4.
[…]
Proof.
See Appendix E-C for the proof.
Theorem 4 quantitatively links differential privacy to algorithmic stability by bounding the mutual information $I[S;WB^{\prime}]$ between the training data and the algorithm’s output. This bound is uniform and scales explicitly with dataset size $n$, alphabet size $|\mathcal{Z}|$, and privacy parameters $(\varepsilon,\delta)$. By translating $(\varepsilon,\delta)$-DP guarantees into a provable stability bound, the theorem establishes a direct connection between privacy and stability-based generalization controls.
Furthermore, Theorem 4 strictly generalizes [23, Proposition 2]: by taking a trivial quantum system ($\dim(B^{\prime})=1$) and setting $\delta=0$, the overhead $h_{\left|\mathcal{Z}\right|}$ vanishes, recovering the classical stability bound $(|\mathcal{Z}|-1)\ln(ne\varepsilon)$.
Remark 3.
The upper-bound obtained in Theorem 4 is independent of $P_{S}$ and thus, Theorem 4 implies that if a quantum learning algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$ satisfies Definition 6, then $\mathcal{N}$ is $\left((|\mathcal{Z}|-1)\ln\left({ne\varepsilon}\right)+\right$. $\left.h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)\right)$-stable (see Definition 1). A similar observation also follows for Corollaries 3 and 4.
In Appendix F-A, we present a detailed comparison of Theorem 4 with [20, Proposition 10]. Additionally, in Appendix F-B, we provide an in-depth comparison of Theorem 4 with the results of [8, Appendix C.7] in the setting of untrusted Data Processors, a topic we will elaborate on in the subsequent section.
IV-C From Stability to Generalization: DP Generalization Guarantees
We now formally demonstrate that differential privacy guarantees generalization. By combining Theorem 4 with Theorem 1 (assuming a trivial system $\bm{\mathfrak{Te}}$), which bounds the expected generalization error via the square root of mutual information, we establish a direct link. Specifically, a $1$-neighbor $(\varepsilon,\delta)$-DP support consistent algorithm limits dependence on individual data points, thereby ensuring robust generalization. We formalize this result in the corollary below.
Corollary 5.
Consider a quantum learning algorithm $\mathcal{N}$ that is $1$-neighbor $(\varepsilon,\delta)$-DP support consistent (see Definition 6) with $\varepsilon\in[\frac{1}{n},1)$ and satisfies condition (12). If the loss operator satisfies (6) (or equivalently (7)) mentioned in Definition 5, for some $\alpha\in(0,\infty)$, then, $\overline{\text{{gen}}}_{\rho}(\mathcal{N})\leq\sqrt{2\alpha^{2}\mathcal{I}_{bound}},$ where $\mathcal{I}_{bound}=\left[(|\mathcal{Z}|-1)\ln\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)\right]$ is the upper bound on the mutual information derived in Theorem 4.
V Untrusted Data Processor and Information Theoretic Admissibility (ITA)
V-A Untrusted Data Processor Model
In the previous sections, we assumed a trusted Data Processor model where the Data Processor reliably executes the privacy-preserving algorithm and releases only the privatized output. We now relax this assumption to address the Untrusted Data Processor scenario. Here, the Data Processor is considered adversarial and may attempt to leak or extract information about the training data $s$ beyond what is contained in the intended output system $B$. To address this privacy threat rigorously, we adopt a worst-case security model where the Data Processor and the Investigator collude or effectively act as a single adversarial entity.
In this setting, the Respondent does not grant the Data Processor access to the raw classical data $s$ directly. Instead, the Respondent provides access only to the set of encoded quantum states $\{\rho_{s}\}_{s}$. The Processor is tasked with running a learning algorithm to produce a hypothesis $w$. Since the Data Processor is untrusted, the learning algorithm is modeled as a single, fixed completely positive, trace-preserving (CP-TP) map $\mathcal{N}$ that must be independent of the specific input index $s$. The total state generated at the output of this process is, $\mathcal{N}(\rho_{s}):=\sum_{w\in\mathcal{W}}\mathcal{N}_{w}(\rho_{s})\otimes|w\rangle\langle w|,$ where each $\mathcal{N}_{w}$ is a completely positive trace non-increasing map summing to $\mathcal{N}$. To ensure privacy, the Respondent mandates that this map $\mathcal{N}$ must satisfy differential privacy constraints with respect to neighboring inputs. The formal definition of such a $1$-neighbor $(\varepsilon,\delta)$-DP support-consistent learning algorithm for an untrusted Data Processor is provided in Definition 11 of Appendix G.
V-B Information-Theoretic Admissibility (ITA): Motivation and Definition
While the Respondent prescribes a specific privacy-preserving algorithm $\mathcal{N}$, an adversarial Data Processor possessing the raw quantum inputs $\{\rho_{s}\}_{s}$ is not technically bound to execute $\mathcal{N}$. The Data Processor aims to extract the maximum possible information about $s$. Therefore, there exists a risk that the Data Processor might execute a strictly more informative algorithm $\mathcal{N}^{\prime}$ and then perform classical or quantum post-processing to simulate the output statistics of the prescribed algorithm $\mathcal{N}$, i.e., they can artificially stitch up the noise after performing a non-private learning algorithm.
If such a scenario is possible, the privacy guarantees calculated based on $\mathcal{N}$ (as mentioned in Definition 11) are rendered void, as the Data Processor effectively holds the information content of $\mathcal{N}^{\prime}$. To formalize this, we introduce the concept of ordering of informativeness between learning algorithms (Definition 12, Appendix G). We say that an algorithm $\mathcal{N}^{\prime}$ is more informative than $\mathcal{N}$ with respect to the ensemble $\{\rho_{s}\}_{s}$ if there exists a post-processing CP-TP map $\Gamma$ (a "simulation map") such that,
$\mathcal{N}(\rho_{s})=\Gamma\circ\mathcal{N}^{\prime}(\rho_{s}),\quad\forall s\in\mathcal{S}.$ (14)
If such a relation holds, the data-processing inequality [21] implies that the mutual information between the input and the output of $\mathcal{N}^{\prime}$ is strictly greater than or equal to that of $\mathcal{N}$.
To certify that a prescribed algorithm $\mathcal{N}$ dominates over every other algorithm $\mathcal{N}^{\prime}$ with respect to the collection $\{\rho_{s}\}_{s}$, we introduce the concept of Information-Theoretic Admissibility (ITA).
Definition 7 (Information-Theoretic Admissibility).
A learning algorithm $\mathcal{N}$ is ITA with respect to the set $\{\rho_{s}\}_{s}$ if there exists no other algorithm $\mathcal{N}^{\prime}$ that is strictly more informative than $\mathcal{N}$, i.e., there does not exist a $\Gamma$ which satisfies (14) for $\mathcal{N}^{\prime}$.
Essentially, if an algorithm is ITA, it implies that the Data Processor is already performing the optimal information extraction allowed by the quantum mechanics formalism.
V-C Quantum Advantage: Privacy under ITA
The imposition of the ITA condition reveals a fundamental divergence between classical and quantum privacy capabilities. We demonstrate that while classical ITA algorithms necessitate a total loss of privacy, quantum mechanics allows for algorithms that are both ITA and privacy-preserving.
The Collapse of Classical Privacy: In the classical domain, the encoded states $\rho_{s}$ effectively behave as probability distributions (or commutative states). In such a scenario, the lemma below shows that simultaneous ITA (optimality) and Differential Privacy (indistinguishability) are impossible.
Lemma 1.
Assume that all states $\{\rho_{s}\}_{s}$ commute (i.e., the classical setting). If there exists no reconstruction map $\Gamma$ such that $\Gamma\circ\mathcal{N}(\rho_{s})=\rho_{s}$, then the algorithm $\mathcal{N}$ is not ITA.
Proof.
See Appendix I
The implication of Lemma 1 is severe: classical ITA algorithms permit full reconstruction of the raw data. Since under ITA the Data Processor effectively holds the raw data, the output-based guarantees of Definition 11 are insufficient. Therefore, for classical ITA algorithms, the privacy condition in Definition 11 must be strengthened to require indistinguishability on the raw states $\rho_{s}$ directly, effectively substituting $\mathcal{N}(\rho_{s})$ with $\rho_{s}$. We discuss more on this in Appendix G-A.
Quantum Privacy via Non-Commutativity: In the quantum regime, the raw inputs $\{\rho_{s}\}_{s}$ may be non-commuting. Quantum mechanics dictates that non-orthogonal states cannot be perfectly distinguished, even by the optimal measurement. Therefore, an algorithm can be ITA—extracting the maximum physically permissible information—without yielding enough information to identify $s$ perfectly.
This allows for the existence of algorithms where the Data Processor performs the best possible measurement (ITA) but remains fundamentally limited by quantum uncertainty, thereby preserving privacy. We illustrate this with a concrete construction:
Example 5 (Quantum ITA algorithm).
Consider the states $\rho_{z}=|\phi_{z,p}\rangle\langle\phi_{z,p}|$ with $|\phi_{z,p}\rangle=\sqrt{1-p}$ $|0\rangle$ $(-1)^{z}\sqrt{p}|1\rangle$. For a dataset $s=(z_{1},\dots,z_{n})$, define the encoded state $|e_{s}\rangle:=\bigotimes_{j=1}^{n}|\phi_{z_{j},1/2}\rangle$. Let $P_{k}$ be the projection onto the subspace spanned by $\{|e_{s}\rangle:|s|_{1}=k\}$ (i.e., the subspace of states corresponds to all strings $s$ with Hamming weight $k$). The untrusted Data Processor applies the projective measurement map $\{\mathcal{N}_{k}\}_{k}$ defined by $\mathcal{N}_{k}(\rho):=P_{k}\rho P_{k}$.
In Example 5, the proposed measurement strategy $\{\mathcal{N}_{k}\}_{k}$ satisfies the Information-Theoretic Admissibility (ITA) condition while maintaining the constraints for privacy. We analyze these properties below:
Justification for ITA: The projectors $\{P_{k}\}_{k}$ define the Hamming weight subspaces. In the orthogonal case ($p=1/2$), the input states $\{\rho_{s}\}_{s}$ are eigenstates of these projectors, making the measurement Quantum Non-Demolition (QND) [6] with $\mathcal{N}(\rho_{s})=\rho_{s}$. In the general case ($p\neq 1/2$), while the measurement may disturb the non-orthogonal states, it remains Information-Theoretically Admissible (ITA). This is because the Hamming weight measurement constitutes the optimal extraction of the ensemble’s geometric parameters; due to the permutation invariance of the encoded states, the Hamming weight acts as a sufficient statistic. Any attempt to extract strictly more information (e.g., individual bit positions) is physically precluded by qubit non-orthogonality, confirming the absence of any strictly superior map $\mathcal{N}^{\prime}$.
Justification for Privacy: The privacy guarantee stems from the distinction between optimal extraction (ITA) and perfect recovery. For $p\neq 1/2$, the qubit states $|\phi_{z,p}\rangle$ are non-orthogonal. Consequently, even though the algorithm extracts the maximal accessible information (the Hamming weight), the inherent quantum uncertainty defined by the Helstrom bound [16] prevents the Investigator from distinguishing the specific string $s$ within the projected subspace. This establishes a scenario where the algorithm is optimal (ITA) yet fundamentally privacy-preserving. Thus, it is meaningful to consider the security condition in Definition 11 as here in the above example, the ITA algorithm is inducing noise while performing the learning algorithm rather than artificially stitching it after.
V-D Distinction from Degradability:
Finally, it is crucial to distinguish ITA from the concept of quantum channel degradation [27, 7]. Degradability asks whether an algorithm can be simulated for any arbitrary input state. In contrast, ITA only asks whether the algorithm can be simulated on the specific training ensemble $\{\rho_{s}\}_{s}$. An algorithm might be non-degradable (secure in a general sense) but still simulable on the specific subspace occupied by the Respondent’s data. Therefore, privacy certification in the untrusted regime must be data-dependent, verifying admissibility explicitly against the geometry of the Respondent’s encoded states.
VI Conclusion
We established an information-theoretic framework for quantum generalization, demonstrating that limited information leakage controls expected generalization error (Theorem 1). Going beyond expected error bounds, we derived a bound on generalization error in probability (Theorem 2) via Sandwiched Rényi divergence and a complementary lower bound on true loss (Theorem 3), effectively sandwiching the risk under a newly introduced Classical-Quantum Sub-Gaussianity (Definition 5).
We further established $(\varepsilon,\delta)$-QDP as a sufficient condition for generalization by deriving a mechanism-agnostic stability bound on the Holevo information (Theorem 4) with logarithmic sample scaling by employing a grid-covering optimization to rigorously handle approximate privacy. Finally, via Information-Theoretic Admissibility (ITA), we demonstrated a fundamental quantum advantage: unlike the classical regime, quantum mechanics permits admissible algorithms that maintain strict privacy against untrusted Data Processors.
Appendix A Proof of Theorem 1
Given the fact that $I[S\bm{\mathfrak{Te}};WB^{\prime}]=D(\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})$, we can lower-bound $D(\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}$ $\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})$ as follows,
$~~~D(\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})$
$\overset{(a)}{\geq}\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\ln\mathrm{Tr}\left[e^{\lambda L^{S\bm{\mathfrak{Te}}WB}}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]$
$=\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]$
$\hskip 50.0pt-\ln\mathrm{Tr}\left[e^{\lambda\left(L^{S\bm{\mathfrak{Te}}WB}-\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]\mathbb{I}^{S\bm{\mathfrak{Te}}WB}\right)}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]$
$\overset{(b)}{\geq}\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]-\frac{\lambda^{2}\alpha^{2}}{2},$ (15)
where $(a)$ follows from the variational lower-bound for the quantum relative entropy (see [15, Theorem $5.9$]) and $(b)$ follows from (7) of Definition 5. Further, we can rewrite (15) as follows,
$\frac{\lambda^{2}\alpha^{2}}{2}-\lambda\left(\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]\right)$
$\hskip 250.0pt+D(\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\geq 0,$
Since the above inequality is a non-negative quadratic equation in $\lambda$ with the coefficient $\frac{\alpha^{2}}{2}\geq 0,$ its discriminant must be non-positive. Hence,
$\left(\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]\right)^{2}\hskip 100.0pt$
$\leq 4\cdot\frac{\alpha^{2}}{2}\cdot D(\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}).$ (16)
Thus, we have,
$\left|\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]\right|\leq\sqrt{2\alpha^{2}I[S\bm{\mathfrak{Te}};WB^{\prime}]}.$ (17)
Therefore, the combination of Definition 4 and (17) yields the following,
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ $\leq\sqrt{2\alpha^{2}I[S\bm{\mathfrak{Te}};WB^{\prime}]}.$ (18)
This completes the proof of Theorem 1.
Appendix B Lower Bound on Expected True Loss of a Quantum Learning Algorithm
In this section, we establish a relationship between the expected true loss $L_{\rho}(\mathcal{N})$ (see Definition 3) and the expected empirical loss $\hat{L}_{\rho}(\mathcal{N})$ (see Definition 2). This bound utilizes the Sandwiched Rényi divergence [19] of order $\gamma\in(1,\infty)$, defined for two quantum states $\rho$ and $\sigma$ as,
$\tilde{D}_{\gamma}(\rho\|\sigma):=\begin{cases}\frac{1}{\gamma-1}\log\mathrm{Tr}\left[\left(\sigma^{\frac{1-\gamma}{2\gamma}}\rho\sigma^{\frac{1-\gamma}{2\gamma}}\right)^{\gamma}\right],&{\mbox{if }}(\rho\ll\sigma),\\ +\infty,&\mbox{else}.\end{cases}$ (19)
B-A Proof of Theorem 3
We denote the product state $\sigma_{\text{prod}}:=\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}$. Our goal is to upper bound the empirical loss $\hat{L}_{\rho}(\mathcal{N})=\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB^{\prime}}\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}]$ in terms of the true loss and the divergence. We begin by expanding the trace using the identity $\mathbb{I}=\sigma_{\text{prod}}^{\frac{\gamma-1}{2\gamma}}\sigma_{\text{prod}}^{\frac{1-\gamma}{2\gamma}}$ and applying a series of information-theoretic inequalities.
Consider the following derivation,
$\hat{L}_{\rho}(\mathcal{N})$ $=\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB^{\prime}}\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}]$
$=\mathrm{Tr}\left[\left(\sigma_{\text{prod}}^{\frac{\gamma-1}{2\gamma}}L^{S\bm{\mathfrak{Te}}WB^{\prime}}\sigma_{\text{prod}}^{\frac{\gamma-1}{2\gamma}}\right)\left(\sigma_{\text{prod}}^{\frac{1-\gamma}{2\gamma}}\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\sigma_{\text{prod}}^{\frac{1-\gamma}{2\gamma}}\right)\right]$
$\overset{(a)}{\leq}\mathrm{Tr}\left[\left|\sigma_{\text{prod}}^{\frac{\gamma-1}{2\gamma}}L^{S\bm{\mathfrak{Te}}WB^{\prime}}\sigma_{\text{prod}}^{\frac{\gamma-1}{2\gamma}}\right|^{\frac{\gamma}{\gamma-1}}\right]^{\frac{\gamma-1}{\gamma}}\mathrm{Tr}\left[\left|\sigma_{\text{prod}}^{\frac{1-\gamma}{2\gamma}}\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\sigma_{\text{prod}}^{\frac{1-\gamma}{2\gamma}}\right|^{\gamma}\right]^{\frac{1}{\gamma}}$
$\overset{(b)}{\leq}\mathrm{Tr}\left[\left(L^{S\bm{\mathfrak{Te}}WB^{\prime}}\right)^{\frac{\gamma}{\gamma-1}}\sigma_{\text{prod}}\right]^{\frac{\gamma-1}{\gamma}}\exp\left(\frac{\gamma-1}{\gamma}\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma_{\text{prod}})\right)$
$\overset{(c)}{\leq}\mathrm{Tr}\left[e^{\frac{\gamma}{\gamma-1}L^{S\bm{\mathfrak{Te}}WB^{\prime}}}\sigma_{\text{prod}}\right]^{\frac{\gamma-1}{\gamma}}\exp\left(\frac{\gamma-1}{\gamma}\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma_{\text{prod}})\right)$
$=\mathrm{Tr}\left[e^{\frac{\gamma}{\gamma-1}\left(L^{S\bm{\mathfrak{Te}}WB^{\prime}}-L_{\rho}(\mathcal{N})\mathbb{I}+L_{\rho}(\mathcal{N})\mathbb{I}\right)}\sigma_{\text{prod}}\right]^{\frac{\gamma-1}{\gamma}}\exp\left(\frac{\gamma-1}{\gamma}\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma_{\text{prod}})\right)$
$=\mathrm{Tr}\left[e^{\frac{\gamma}{\gamma-1}\left(L^{S\bm{\mathfrak{Te}}WB^{\prime}}-L_{\rho}(\mathcal{N})\mathbb{I}\right)}e^{\frac{\gamma}{\gamma-1}L_{\rho}(\mathcal{N})\mathbb{I}}\sigma_{\text{prod}}\right]^{\frac{\gamma-1}{\gamma}}\exp\left(\frac{\gamma-1}{\gamma}\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma_{\text{prod}})\right)$
$=\mathrm{Tr}\left[e^{\frac{\gamma}{\gamma-1}\left(L^{S\bm{\mathfrak{Te}}WB^{\prime}}-L_{\rho}(\mathcal{N})\mathbb{I}\right)}\sigma_{\text{prod}}\right]^{\frac{\gamma-1}{\gamma}}\exp\left(L_{\rho}(\mathcal{N})+\frac{\gamma-1}{\gamma}\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma_{\text{prod}})\right)$
$\overset{(d)}{\leq}\exp\left(\frac{\gamma\alpha^{2}}{2(\gamma-1)}+L_{\rho}(\mathcal{N})+\frac{\gamma-1}{\gamma}\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma_{\text{prod}})\right).$
Justification of inequalities:
  • $(a)$
    Follows from the non-commutative Hölder inequality [3],
    $\mathrm{Tr}[|AB|]\leq\left(\mathrm{Tr}[A^{p}]\right)^{1/p}\left(\mathrm{Tr}[B^{q}]\right)^{1/q},$ (20)
    for any two positive opertors $A$ and $B$, where we identify the operators $A=\sigma_{\text{prod}}^{\frac{\gamma-1}{2\gamma}}L^{S\bm{\mathfrak{Te}}WB^{\prime}}\sigma_{\text{prod}}^{\frac{\gamma-1}{2\gamma}}$ and $B=\sigma_{\text{prod}}^{\frac{1-\gamma}{2\gamma}}\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\sigma_{\text{prod}}^{\frac{1-\gamma}{2\gamma}}$ to be positive operators, with conjugate exponents $p=\frac{\gamma}{\gamma-1}$ and $q=\gamma$.
  • $(b)$
    The first factor follows from the Araki-Lieb-Thirring inequality [1, 17] $\mathrm{Tr}[(BAB)^{r}]\leq\mathrm{Tr}[A^{r}B^{2r}]$ (with $B=\sigma_{\text{prod}}^{\frac{\gamma-1}{2\gamma}}$, $A=L^{S\bm{\mathfrak{Te}}WB^{\prime}}$, and $r=\frac{\gamma}{\gamma-1}$). The second factor follows directly from the definition of the Sandwiched Rényi divergence $\tilde{D}_{\gamma}$.
  • $(c)$
    Follows from the operator inequality $X^{p}\leq e^{pX}$ for any positive operator $X\geq 0$ and $p>0$. Here, we apply this to the operator $X=L^{S\bm{\mathfrak{Te}}WB^{\prime}}$ with $p=\frac{\gamma}{\gamma-1}$.
  • $(d)$
    Follows from the Classical-Quantum Sub-Gaussian assumption (Definition 5). By setting $\lambda=\frac{\gamma}{\gamma-1}$, the assumption guarantees $\mathrm{Tr}\left[e^{\lambda\left(L^{S\bm{\mathfrak{Te}}WB^{\prime}}-L_{\rho}(\mathcal{N})\mathbb{I}\right)}\sigma_{\text{prod}}\right]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2}}$. Raising this to the power of $\frac{1}{\lambda}=\frac{\gamma-1}{\gamma}$ yields the term $\exp\left(\frac{\lambda\alpha^{2}}{2}\right)=\exp\left(\frac{\gamma\alpha^{2}}{2(\gamma-1)}\right)$.
Rearranging the terms in the final inequality to lower-bound $\exp(L_{\rho}(\mathcal{N}))$ yields the statement of the theorem.
Further, if the loss observables $\{L(w,s)\}$ are strictly bounded between 0 and $\mathbb{I}$, we can derive a tighter multiplicative lower bound on the expected true loss that does not depend on the sub-Gaussian parameter $\alpha$.
Corollary 6.
Let $\mathcal{N}$ be a quantum learning algorithm. Assume the loss operators are bounded such that $0\preceq L(w,s)\preceq\mathbb{I}$, for all $w,s$. For any sandwiched Rényi divergence order $\gamma>1$, the expected true loss is lower bounded by the empirical loss as follows,
$L_{\rho}(\mathcal{N})\geq\hat{L}^{\frac{\gamma}{\gamma-1}}_{\rho}(\mathcal{N})\exp\left(-\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right).$ (21)
Remark 4.
Corollary 6 can be viewed as a quantum analogue of [18, Theorem 3]. In the classical setting, [18, Theorem 3] derived a similar lower bound on the true risk in terms of the empirical risk and the classical Rényi divergence under the assumption of bounded loss functions. Our result extends this bound to the quantum learning framework, where the non-commutativity of the state and loss operators necessitates the use of the Sandwiched Rényi divergence $\tilde{D}_{\gamma}$ (which reduces to the classical Rényi divergence when states commute) and the utilization of the non-commutative Hölder inequality to separate the statistical fluctuations from the dependency structure.
Appendix C Probabilistic Bound on Generalization Error via Sandwiched Rényi divergence
In this section, we establish an upper-bound on the absolute generalization error in probability using the Sandwiched Rényi divergence. In contrast to the bound on the expected generalization error (Theorem 1), which is based on standard Mutual Information and only controls the error on average, our focus here is on obtaining guarantees that hold with high probability (confidence level $1-\delta$). Such guarantees are essential in safety-critical settings, where average-case performance is not enough; one must ensure the error remains small with high confidence.
Achieving this type of high-probability guarantee requires controlling higher-order moments of the dependence between the data and the learned hypothesis. This dependence is quantified by the Sandwiched Rényi divergence. Even though the sandwiched Rényi divergence ($\tilde{D}_{\gamma}$) gives a larger value than standard Mutual Information for $\gamma\in(1,\infty)$, using this stronger measure allows us to separate the randomness of the loss function from how much the algorithm relies on the training data with the help of non-commutative Hölder’s inequality [3]. This separation is necessary to guarantee that the model performs well even in the worst-case scenarios.
C-A Definitions and Assumptions for Probabilistic Bounds
In the general setting discussed previously, the generalization error bound relies on the mutual information $I(S;W)$, which captures the aggregate dependency between the dataset and the hypothesis. However, to derive probabilistic upper-bounds that hold with high probability and decay exponentially with the sample size nn—it is standard in statistical learning theory to assume structural independence in the data generation and processing [13, 18].
Without such assumptions, worst-case correlations between data points could prevent the empirical loss from concentrating around its true mean. Therefore, we adopt an independent and identically distributed (i.i.d.) framework. Physically, this corresponds to a scenario where the Data Processor processes each incoming quantum data state independently (e.g., via parallel quantum channels or distinct experimental repetitions) before aggregating the results to form a hypothesis.
I.I.D. Structure of Data and Algorithm:
We assume the dataset $S=\{Z_{1},\dots,Z_{n}\}$ consists of $n$ i.i.d. random variables, where each $Z_{i}\sim P_{Z}$. Commensurate with this, we assume the quantum learning algorithm $\mathcal{N}$ respects this independence by acting on each data encoding locally. Specifically, the global channel decomposes as a tensor product,
$\mathcal{N}^{(S)}_{w}:=\bigotimes_{i=1}^{n}\mathcal{N}^{(Z_{i})}_{w}$,
where each local map $\mathcal{N}^{(Z_{i})}_{w}:\mathcal{H}^{\hat{\bm{\mathfrak{Tr}}}}\to\mathcal{H}^{\hat{B}}$ acts on the input state $\rho_{Z_{i}}$ particular to the $i$-th datapoint. Consequently, the residual quantum output system decomposes as $B^{\prime}:=\hat{B}^{\otimes n}$.
Decomposition of Loss:
Consistent with the independence of the processing, we assume the global loss observable $L(w,s)$ is the average of local loss observables acting on the individual subsystems. This decomposition is crucial because it allows us to view the total generalization error as the average of $n$ independent random variables, thereby enabling the use of Chernoff-type bounds where the variance scales inversely with $n$.
$L(w,s):=\frac{1}{n}\sum_{i=1}^{n}(\mathbb{I}^{\hat{\bm{\mathfrak{Te}}}}\otimes\mathbb{I}^{\hat{B}})^{\otimes(i-1)}\otimes\hat{L}(w,z_{i})\otimes(\mathbb{I}^{\hat{\bm{\mathfrak{Te}}}}\otimes\mathbb{I}^{\hat{B}})^{\otimes(n-i)},$ (22)
where $\hat{L}(w,z_{i})$ is the local loss observable for the $i$-th data point.
To guarantee exponential probability, we require the local loss operators to satisfy a sub-Gaussian condition. This is a standard regularity condition ensuring the tails of the loss distribution decay sufficiently fast.
Definition 8 (Classical-Quantum Local $\alpha$-Sub-Gaussianity).
For a fixed parameter $\alpha\in(0,\infty),$ the collection $\{\hat{L}(w,z)\}$ of local loss operators is said to be an $\alpha$-sub-Gaussian collection if, for every $\lambda\in\mathbb{R}$, the centered local moment generating function satisfies,
$\mathbb{E}\left[\mathrm{Tr}\left[e^{\lambda\left(\hat{L}(Z_{i},W)-\mathbb{E}\left[\mathrm{Tr}\left[\hat{L}(Z_{i},W)\left(\rho^{\bm{\mathfrak{Te}}}_{Z_{i}}\otimes(\sigma^{\mathcal{N}}_{W})^{\hat{B}}\right)\right]\mathbb{I}^{\hat{\bm{\mathfrak{Te}}}\hat{B}}\right]\right)}\left(\rho^{\hat{\bm{\mathfrak{Te}}}}_{Z_{i}}\otimes(\sigma^{\mathcal{N}}_{W})^{\hat{B}}\right)\right]\right]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2}},$ (23)
where the expectations are taken with respect to the product distribution $P_{Z}\times{P}^{\mathcal{N}}_{{W}}$.
Remark 5 (Scaling of the Global Variance).
It is important to highlight that the local assumption (23) implies a strictly tighter bound for the global loss operator. Specifically, due to the tensor product structure and the independence of $Z_{i}$, the global condition holds with a variance proxy that scales as $1/n$,
$\mathrm{Tr}\left[e^{\lambda\left(L^{S\bm{\mathfrak{Te}}WB}-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\mathbb{I}^{S\bm{\mathfrak{Te}}WB}\right)}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2n}}.$ (24)
This scaling is a direct consequence of the additivity of cumulants for independent variables (or technically, via Jensen’s inequality for the operator exponential). This $1/n$ factor is precisely what allows the generalization bound to vanish as the dataset size increases.
A direct consequence of this scaling is that the expected generalization error bound from Theorem 1 also benefits from the sample size. Substituting the variance proxy $\frac{\alpha^{2}}{n}$ into the framework of Theorem 1 yields us the following result
Corollary 7 (Expected Generalization Bound under I.I.D. Assumption).
For a fixed $\alpha\in(0,\infty),$ if the loss operators for a quantum learning algorithm $\mathcal{N}$, satisfy Definition 8, then, we have,
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})\leq\sqrt{\frac{2\alpha^{2}}{n}I[S\bm{\mathfrak{Te}};WB^{\prime}]}.$
This explicitly demonstrates the $\sqrt{1/n}$ convergence rate for the expected error, confirming that algorithmic stability (bounded mutual information) leads to vanishing generalization error as $n\to\infty$.
C-B Proof of Theorem 2
Under this setting mentioned in Subsection C-A, we formally define the random variable representing the absolute deviation of the generalization error.
Definition 9 (Absolute Generalization Error Deviation).
Let $\mathcal{N}$ be a quantum learning algorithm with the i.i.d. structure defined above. Using [26, Definition 2020], for a given $w\in\mathcal{W}$, we define the generalization error random variable as the absolute difference between the empirical loss and the true loss $L_{\rho}(\mathcal{N},w)$, (see [26, Definition $17$]),
$~~~\textnormal{gen}_{\rho}(\mathcal{N},S,w)$
$:=\left|\mathrm{Tr}\left[L(S,w)(\sigma^{\mathcal{N}}_{S,w})^{\bm{\mathfrak{Te}}B^{\prime}}\right]-L_{\rho}(\mathcal{N},w)\right|$
[…]
To get a quantum version of the probabilistic bound obtained in [13, Corollary $2$], we make a quantum version of the sub-Gaussianity assumption on the loss functions mentioned in Corollary $2$ on [13, Page $8$].
Definition 10 (Conditional Classical-Quantum Local $\alpha$-Sub-Gaussianity).
For a fixed parameter $\alpha\in(0,\infty),$ for every $w\in\mathcal{W}$, the collection $\{\hat{L}(w,z)\}_{z}$ of local loss operators is said to be an $\alpha$-sub-Gaussian collection if, for every $\lambda\in\mathbb{R}$, the centered local moment generating function satisfies,
$\mathbb{E}\left[\mathrm{Tr}\left[e^{\lambda\left(\hat{L}(Z,w)-\mathbb{E}\left[\mathrm{Tr}\left[\hat{L}(Z,w)\left(\rho^{\bm{\mathfrak{Te}}}_{Z}\otimes(\sigma^{\mathcal{N}}_{w})^{\hat{B}}\right)\right]\mathbb{I}^{\hat{\bm{\mathfrak{Te}}}\hat{B}}\right]\right)}\left(\rho^{\hat{\bm{\mathfrak{Te}}}}_{Z}\otimes(\sigma^{\mathcal{N}}_{w})^{\hat{B}}\right)\right]\right]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2}},$ (25)
where the expectations are taken with respect to the distribution $P_{Z}$. Note that in the special case, there is no quantum system, then for every $w$, (25) reduces to the sub-Gaussianity assumption on the loss functions mentioned in Corollary $2$ on [13, Page $8$].
It is important to highlight that the local assumption (25) is equivalent to the following condition,
$\mathrm{Tr}\left[e^{\lambda\left(L^{S\bm{\mathfrak{Te}}B}_{w}-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}B^{\prime}}_{w}(\sigma^{S\bm{\mathfrak{Te}}}\otimes(\sigma^{\mathcal{N}}_{w})^{B^{\prime}})\right]\mathbb{I}^{S\bm{\mathfrak{Te}}B}\right)}(\sigma^{S\bm{\mathfrak{Te}}}\otimes(\sigma^{\mathcal{N}}_{w})^{B^{\prime}})\right]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2n}}.$ (26)
where, for every $w\in\mathcal{W}$, we define, $L^{S\bm{\mathfrak{Te}}B^{\prime}}_{w}:=\sum_{s\in\mathcal{S}}\overset{S}{|s\rangle\langle s|}\otimes\overset{\bm{\mathfrak{Te}}B^{\prime}}{L(s,w)}.$
We now prove Theorem 2 by defining $\mathcal{E}_{S,W}:=\mathrm{Tr}\left[L(S,W)(\sigma^{\mathcal{N}}_{S,W})^{\bm{\mathfrak{Te}}B^{\prime}}\right]-L_{\rho}(\mathcal{N},W)$. We are interested in bounding $\Pr\{|\mathcal{E}_{S,W}|>\varepsilon\}$. By the union bound, this probability is written as,
$\Pr\{|\mathcal{E}_{S,W}|>\varepsilon\}=\Pr\{\mathcal{E}_{S,W}>\varepsilon\}+\Pr\{\mathcal{E}_{S,W}<-\varepsilon\}=\Pr\{\mathcal{E}_{S,W}>\varepsilon\}+\Pr\{-\mathcal{E}_{S,W}>\varepsilon\}.$
We first bound the positive deviation $\Pr\{\mathcal{E}_{S,W}>\varepsilon\}$. Applying the Markov inequality for any $\lambda>0$, we have,
$~~~\Pr\{\mathcal{E}_{S,W}>\varepsilon\}$
$=\Pr_{(S,W)\sim P^{\mathcal{N}}_{SW}}\left\{\mathrm{Tr}\left[(L(S,W)-L_{\rho}(\mathcal{N},W)\mathbb{I})(\sigma^{\mathcal{N}}_{S,W})^{\bm{\mathfrak{Te}}B^{\prime}}\right]>\varepsilon\right\}$
$=E_{W\sim P^{\mathcal{N}}_{W}}\left[\Pr_{S\sim P^{\mathcal{N}}_{S|W}}\left\{\mathrm{Tr}\left[(L(S,W)-L_{\rho}(\mathcal{N},W)\mathbb{I})(\sigma^{\mathcal{N}}_{S,W})^{\bm{\mathfrak{Te}}B^{\prime}}\right]>\varepsilon\right\}\right].$ (27)
For a fixed $w\in\mathcal{W}$, we have,
$~~~\Pr_{S\sim P^{\mathcal{N}}_{S|W=w}}\left\{\mathrm{Tr}\left[(L(S,w)-L_{\rho}(\mathcal{N},w)\mathbb{I})(\sigma^{\mathcal{N}}_{S,w})^{\bm{\mathfrak{Te}}B^{\prime}}\right]>\varepsilon\right\}$
$\overset{(a)}{\leq}\Pr_{S\sim P^{\mathcal{N}}_{S|W=w}}\left\{\mathrm{Tr}\left[e^{\lambda(L(S,w)-L_{\rho}(\mathcal{N},w)\mathbb{I})}(\sigma^{\mathcal{N}}_{S,w})^{\bm{\mathfrak{Te}}B^{\prime}}\right]>e^{\lambda\varepsilon}\right\}$
$\leq e^{-\lambda\varepsilon}\mathbb{E}_{S|W=w}\left[\mathrm{Tr}\left[e^{\lambda(L(S,w)-L_{\rho}(\mathcal{N},w)\mathbb{I})}(\sigma^{\mathcal{N}}_{S,w})^{\bm{\mathfrak{Te}}B^{\prime}}\right]\right]$
$=e^{-\lambda\varepsilon}\mathrm{Tr}\left[e^{\lambda(L^{S\bm{\mathfrak{Te}}B^{\prime}}_{w}-L_{\rho}(\mathcal{N},w)\mathbb{I})}\sigma^{S\bm{\mathfrak{Te}}B^{\prime}}_{\mathcal{N},w}\right].$ (28)
where ($a$) follows from the fact that for any Hermitian operator $H$ and density matrix $\rho$, the convexity bound $\mathrm{Tr}[e^{H}\rho]\geq e^{\mathrm{Tr}[H\rho]}$ holds and in (28), we define $\sigma^{S\bm{\mathfrak{Te}}B^{\prime}}_{\mathcal{N},w}:=\sum_{s\in\mathcal{S}}P^{\mathcal{N}}_{S|W=w}(s)$ $|s\rangle\langle s|\otimes(\sigma^{\mathcal{N}}_{s,w})^{\bm{\mathfrak{Te}}B^{\prime}}$.
To bound the trace term in Eq. (28), we introduce the product state $\sigma_{\text{prod},w}:=\sigma^{S\bm{\mathfrak{Te}}}\otimes(\sigma^{\mathcal{N}}_{w})^{B^{\prime}}$. Thus, we have,
$~~~\mathrm{Tr}\left[e^{\lambda(L^{S\bm{\mathfrak{Te}}B^{\prime}}_{w}-L_{\rho}(\mathcal{N},w)\mathbb{I})}\sigma^{S\bm{\mathfrak{Te}}B^{\prime}}_{\mathcal{N},w}\right]$
$=\mathrm{Tr}\left[\left(\sigma_{\text{prod},w}^{\frac{\gamma-1}{2\gamma}}e^{\lambda(L^{S\bm{\mathfrak{Te}}B^{\prime}}_{w}-L_{\rho}(\mathcal{N},w)\mathbb{I})}\sigma_{\text{prod},w}^{\frac{\gamma-1}{2\gamma}}\right)\left(\sigma_{\text{prod},w}^{\frac{1-\gamma}{2\gamma}}\sigma^{S\bm{\mathfrak{Te}}B^{\prime}}_{\mathcal{N},w}\sigma_{\text{prod},w}^{\frac{1-\gamma}{2\gamma}}\right)\right].$ (29)
We invoke the non-commutative Hölder’s inequality,
$\left|\mathrm{Tr}[AB]\right|\leq\left(\mathrm{Tr}[\left|A\right|^{p}]\right)^{\frac{1}{p}}\left(\mathrm{Tr}[\left|B\right|^{q}]\right)^{\frac{1}{q}},$
choosing the conjugate exponents $q=\gamma$ and $p=\frac{\gamma}{\gamma-1}$, and defining the operators
$A:=\sigma_{\text{prod},w}^{\frac{\gamma-1}{2\gamma}}e^{\lambda(L^{S\bm{\mathfrak{Te}}B^{\prime}}_{w}-L_{\rho}(\mathcal{N},w)\mathbb{I})}\sigma_{\text{prod},w}^{\frac{\gamma-1}{2\gamma}},\quad\quad B:=\sigma_{\text{prod},w}^{\frac{1-\gamma}{2\gamma}}\sigma^{S\bm{\mathfrak{Te}}B^{\prime}}_{\mathcal{N},w}\sigma_{\text{prod},w}^{\frac{1-\gamma}{2\gamma}}.$
[…]
1. Analysis of Term I (Algorithm’s Data Dependency):
By the definition of sandwiched Rényi divergence and the fact that $B$ is a positive operator, we have,
$\left(\mathrm{Tr}[\left|B\right|^{\gamma}]\right)^{\frac{1}{\gamma}}$ $=\left(\mathrm{Tr}\left[\left(\sigma_{\text{prod},w}^{\frac{1-\gamma}{2\gamma}}\sigma^{S\bm{\mathfrak{Te}}B^{\prime}}_{\mathcal{N},w}\sigma_{\text{prod},w}^{\frac{1-\gamma}{2\gamma}}\right)^{\gamma}\right]\right)^{\frac{1}{\gamma}}=\exp\left(\frac{\gamma-1}{\gamma}\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}B^{\prime}}_{\mathcal{N},w}\|\sigma_{\text{prod},w})\right).$ (30)
2. Analysis of Term II (Randomness of the Loss operators):
The combination of the fact that $B$ is a positive operator and the Araki-Lieb-Thirring inequality ($\mathrm{Tr}[(YXY)^{r}]\leq\mathrm{Tr}[Y^{r}X^{r}Y^{r}]$) with $r=p=\frac{\gamma}{\gamma-1}$, yields the following,
$\left(\mathrm{Tr}[\left|A\right|^{p}]\right)^{\frac{1}{p}}$ $=\left(\mathrm{Tr}\left[\left(\sigma_{\text{prod},w}^{\frac{\gamma-1}{2\gamma}}e^{\lambda(L^{S\bm{\mathfrak{Te}}B^{\prime}}_{w}-L_{\rho}(\mathcal{N},w)\mathbb{I})}\sigma_{\text{prod},w}^{\frac{\gamma-1}{2\gamma}}\right)^{p}\right]\right)^{\frac{1}{p}}$
$\leq\left(\mathrm{Tr}\left[\sigma^{1/2}_{\text{prod}}e^{\frac{\gamma\lambda}{\gamma-1}(L^{S\bm{\mathfrak{Te}}B^{\prime}}_{w}-L_{\rho}(\mathcal{N},w)\mathbb{I})}\sigma^{1/2}_{\text{prod}}\right]\right)^{\frac{\gamma-1}{\gamma}}$
$=\left(\mathrm{Tr}\left[e^{\frac{\gamma\lambda}{\gamma-1}(L^{S\bm{\mathfrak{Te}}B^{\prime}}_{w}-L_{\rho}(\mathcal{N},w)\mathbb{I})}\sigma_{\text{prod},w}\right]\right)^{\frac{\gamma-1}{\gamma}}.$ (31)
Invoking the Classical-Quantum Sub-Gaussian assumption ((26)) for $\alpha$ and setting $\lambda\leftarrow\frac{\gamma\lambda}{\gamma-1}$ in (7), we have,
$\left(\mathrm{Tr}[\left|A\right|^{p}]\right)^{\frac{1}{p}}$ $\leq\left(\exp\left(\frac{1}{2n}\left(\frac{\gamma\lambda}{\gamma-1}\right)^{2}\alpha^{2}\right)\right)^{\frac{\gamma-1}{\gamma}}=\exp\left(\frac{\gamma\lambda^{2}\alpha^{2}}{2n(\gamma-1)}\right).$ (32)
3. Aggregation and Global Divergence:
Substituting (30) and (32) back into (28), and then computing the expectation over $W$ in (27), yields us,
$~~~\Pr\{\mathcal{E}_{S,W}>\varepsilon\}$
$\leq e^{-\lambda\varepsilon}\exp\left(\frac{\gamma\lambda^{2}\alpha^{2}}{2n(\gamma-1)}\right)\mathbb{E}_{W\sim P^{\mathcal{N}}_{W}}\left[\exp\left(\frac{\gamma-1}{\gamma}\tilde{D}{\gamma}(\sigma^{S\bm{\mathfrak{Te}}B^{\prime}}_{\mathcal{N},w}\|\sigma_{\text{prod},w})\right)\right]$
$\overset{(a)}{\leq}e^{-\lambda\varepsilon}\exp\left(\frac{\gamma\lambda^{2}\alpha^{2}}{2n(\gamma-1)}\right)\left(\mathbb{E}_{W\sim P^{\mathcal{N}}_{W}}\left[\exp\left((\gamma-1)\tilde{D}{\gamma}(\sigma^{S\bm{\mathfrak{Te}}B^{\prime}}_{\mathcal{N},w}\|\sigma_{\text{prod},w})\right)\right]\right)^{\frac{1}{\gamma}}$
$\overset{(b)}{\leq}e^{-\lambda\varepsilon}\exp\left(\frac{\gamma\lambda^{2}\alpha^{2}}{2n(\gamma-1)}\right)\left(\exp\left((\gamma-1)\tilde{D}{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right)\right)^{\frac{1}{\gamma}}$
$=\exp\left(-\lambda\varepsilon+\frac{\gamma\lambda^{2}\alpha^{2}}{2n(\gamma-1)}\right)\exp\left(\frac{\gamma-1}{\gamma}\tilde{D}{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right)$
$=\exp\left(-\lambda\varepsilon+\frac{\gamma\lambda^{2}\alpha^{2}}{2n(\gamma-1)}+\frac{\gamma-1}{\gamma}\tilde{D}{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right)$ (33)
where $(a)$ follows from Jensen’s inequality and the concavity of the function $f(x)=x^{\frac{1}{\gamma}}$ for $\gamma>1$ and $(b)$ follows since the joint state $\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}$ and the product state $\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}$ are block-diagonal with respect to the classical system $W$, the divergence decomposes as,
$\tilde{D}{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})=\frac{1}{\gamma-1}\ln\mathbb{E}_{W}\left[\exp\left((\gamma-1)\tilde{D}{\gamma}(\sigma^{S\bm{\mathfrak{Te}}B^{\prime}}_{\mathcal{N},w}\|\sigma_{\text{prod},w})\right)\right].$
Minimizing the exponent $f(\lambda)=-\lambda\varepsilon+\frac{\gamma\alpha^{2}}{2n(\gamma-1)}\lambda^{2}$ yields $\lambda^{*}=\frac{n\varepsilon(\gamma-1)}{\gamma\alpha^{2}}$, resulting in,
$\Pr\{\mathcal{E}_{S,W}>\varepsilon\}\leq\exp\left(-\frac{\gamma-1}{\gamma}\left(\frac{n\varepsilon^{2}}{2\alpha^{2}}-\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma_{\text{prod},w})\right)\right).$ (34)
By symmetry of the sub-Gaussian assumption, the same bound holds for the negative deviation $\Pr\{-Z>\varepsilon\}$. Thus, for the absolute deviation $|Z|$, we have,
$\Pr\{|\mathcal{E}_{S,W}|>\varepsilon\}\leq 2\exp\left(-\frac{\gamma-1}{\gamma}\left(\frac{n\varepsilon^{2}}{2\alpha^{2}}-\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma_{\text{prod},w})\right)\right).$ (35)
4. Inversion for High-Probability Guarantee:
We set the upper bound equal to the confidence level $\delta\in(0,1)$,
$\delta=2\exp\left(-\frac{\gamma-1}{\gamma}\left(\frac{n\varepsilon^{2}}{2\alpha^{2}}-\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma_{\text{prod},w})\right)\right).$
Finally, solving for $\varepsilon$, we have,
$\varepsilon=\sqrt{\frac{2\alpha^{2}}{n}\left(\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})+\frac{\gamma}{\gamma-1}\ln\frac{2}{\delta}\right)}.$
This completes the proof of Theorem 2.
Remark 6 (Comparison with Classical Bounds).
In the special case where the quantum input subsystem $\bm{\mathfrak{Te}}$ and the algorithm’s internal quantum output $B^{\prime}$ are trivial (i.e., the systems are purely classical), the term $\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})$ reduces to Sibson’s mutual information [25] $I_{\gamma}(S;W)$ of order $\gamma$. Consequently, the bound derived in Theorem 2 recovers the exact form of Corollary 2 in [13].
Appendix D Comparison of various Generalization Error Bounds obtained in this manuscript with the Prior Works
In this section, we compare the upper-bounds on the generalization error (Theorem 1 and Theorem 2) derived in this work—both in expectation and in probability—with existing results.
D-A Comparision of Theorem 1 with Corollary $23$ of [8]
We contrast our upper-bound on the expected generalization error (Theorem 1) with Corollary $23$ of [8], highlighting the key advantages of our framework.
  • Simplified Sub-Gaussianity Assumptions: In [8, Corollary $23$], the authors impose the following two separate point-wise holding sub-Gaussianity requirements,
    $\mathrm{Tr}\left[e^{\lambda(L(s,w)-f(s,w)\mathbb{I}^{\bm{\mathfrak{Te}}B^{\prime}})}(\rho^{\bm{\mathfrak{Te}}}_{s}\otimes(\sigma^{\mathcal{N}}_{s,w})^{B^{\prime}})\right]$ $\leq e^{\frac{\mu^{2}\lambda^{2}}{2}},\qquad\forall(s,w)\in\mathcal{S}\times\mathcal{W},$ (QMGF)
    $\mathbb{E}_{S\sim P^{m}}\left[e^{\lambda(f(S,w)-\mathbb{E}_{\tilde{S}}[f(\tilde{S},w)])}\right]$ $\leq e^{\frac{\tau^{2}\lambda^{2}}{2}},\qquad\forall w\in\mathcal{W}.$ (CMGF)
    for some fixed $\mu,\tau>0$, where $f(s,w)\coloneqq\mathrm{Tr}[L(s,w)(\rho^{\bm{\mathfrak{Te}}}_{s}\otimes(\sigma^{\mathcal{N}}_{s,w})^{B^{\prime}})]$. However, unlike Eqs. (QMGF) and (CMGF) in [8], Theorem 1 requires only a single condition mentioned in Definition 5. Crucially, our assumption holds in expectation over $P_{S}\times P_{W}$, rather than for worst-case pairs.
  • Unified Information Measure: The upper-bound obtained in [8, Corollary $23$] contains separated classical and quantum information terms because of the separated sub-Gaussianity assumption mentioned in (QMGF) and (CMGF) respectively. In constrast, our bound on $\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ in Theorem 1 relies on a single information-theoretic quantity, which unifies classical and quantum dependencies.
  • Failure of Stability Implications: In scenarios where testing and training data are uncorrelated, the bound in [8, Corollary $23$] reduces to a purely classical term and therefore it will not account for the quantum system $B^{\prime}$. Therefore, from [8, Corollary $23$] it is not possible to show that stability implies generalizability. In contrast, Theorem 1 avoids this limitation, validating the definition of expected true loss proposed in [26] as its correct formulation. A justification for the correctness of (3) is also given in [26].
With this correct definition of true loss, the bound obtained in [8, Corollary 23] translates to [26, Theorem 1]. In Appendix D-B, we make a comparison of Theorem 1 with [26, Theorem 1] for the case when $\alpha=\mu=\tau,$ where $\alpha,\mu$ and $\tau$ denote the sub-Gaussianity parameters appearing in (6), (QMGF), and (CMGF), respectively.
D-B Numerical Comparison of Theorem 1 with [26, Theorem 1]
In this appendix, we numerically validate our theoretical results by comparing them against the bounds established in [26]. We utilize the classical-quantum toy example described in [26, Section VI] to demonstrate the tightness of our mutual information-based approach.
For this comparison, we evaluate the following two quantities under the condition that the sub-Gaussianity parameters satisfy $\mu=\tau=\alpha$.
Figure 3: Numerical comparison of the generalization error bounds for the classical-quantum toy example in [26]. (a) Comparison of $\mathcal{B}_{\text{MI}}$ (36) and $\mathcal{B}_{\text{SEP}}$ (37) as a function of the prior probability $p\in[0.25,0.75]$. (b) Comparison of $\mathcal{B}_{\text{MI}}$ (36) and $\mathcal{B}_{\text{SEP}}$ (37) as a function of the sub-Gaussianity parameter $\alpha\in[0.1,1]$ for a fixed prior $p=0.4$. In both regimes, our bound $\mathcal{B}_{\text{MI}}$ (blue) provides a strictly tighter upper bound than $\mathcal{B}_{\text{SEP}}$ (orange).
  1. 1.
    Our Mutual Information Bound ($\mathcal{B}_{\text{MI}}$): Derived from Theorem 1, this bound relies on the total mutual information between the input and the output system. Due to the independence of the test and train systems conditioned on $Z$ in this example, the term $I(Z,\bm{\mathfrak{Te}};WB^{\prime})$ simplifies, yielding:
    ${\color[rgb]{0.12109375,0.46484375,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.46484375,0.70703125}\mathcal{B}_{\text{MI}}=\sqrt{2\alpha^{2}I(Z;WB^{\prime})}}.$ (36)
  2. 2.
    The Separated Bound from [26] ($\mathcal{B}_{\text{SEP}}$): We compare against the bound in [26, Eq. (88), Theorem 1], which separates the classical and quantum contributions. In this specific toy example, the first term of their theorem vanishes, reducing the bound to:
    ${\color[rgb]{1,0.49609375,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49609375,0.0546875}\mathcal{B}_{\text{SEP}}=\mathbb{E}_{Z,W}\left[\sqrt{2\alpha^{2}D\left(\sigma^{B^{\prime}}_{(Z,W)}\middle\|\sigma^{B^{\prime}}_{W}\right)}\right]+\sqrt{2\alpha^{2}I(Z;W)}}.$ (37)
As detailed in Figure 3 above, for comparison, we plot (36) and (37) for the example mentioned in [26, Subsection-VI].
D-C Comparision of Theorem 1 with Theorem $2$ of [26]
In Theorem $2$ of [26], the authors obtained an upper-bound on expected generalization bounds of quantum learning algorithms, in terms of Rényi divergences. We contrast our upper-bound on the expected generalization error (Theorem 1) with their results, highlighting the key advantages of our framework.
  • Simplified Sub-Gaussianity Assumptions: In [26, Theorem $2$], the authors introduce five distinct point-wise sub-Gaussianity assumptions, as specified in [26, Assumption $5$ and $6$]. These conditions are needed there to establish an upper bound in terms of the Rényi divergence. In contrast, Theorem 1 replaces Assumptions $5$ and $6$ in [26] with a single requirement, given in Definition 5. A key difference is that our condition is formulated in expectation with respect to $P_{S}\times P_{W}$, rather than being imposed for all worst-case pairs.
  • Unified Information Measure: The upper bound derived in [26, Theorem $2$] involves two terms based on quantum Rényi divergence and one term based on classical Rényi divergence. This structure stems from the separate sub-Gaussianity assumptions formulated in [26, Assumption $5$ and $6$]. By contrast, our bound in Theorem 1 is expressed in terms of a single information-theoretic quantity that simultaneously captures both classical and quantum dependencies.
D-D Comparison of Theorem 2 with Theorem 4 of [26]
We contrast our concentration bound (Theorem 2) with Theorem 4 of [26], highlighting three key advantages of our framework.
  • Simplified Sub-Gaussianity Assumptions: The result in [26] necessitates a quantum learning framework with separate pointwise sub-Gaussianity conditions for the quantum posterior states and the classical hypothesis distribution. We simplify these requirements significantly, relying on only a single sub-Gaussian condition (Eq. (23)) on the individual loss operators.
  • Unified Information Measure: The framework in [26] employs a "separated" approach that sums classical mutual information and a distinct quantum divergence term. This decoupling is analytically complex and often results in looser bounds. In contrast, our bound relies on a single global divergence, $\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})$, which captures classical and quantum dependencies jointly within a unified measure.
  • Average-Case vs. Worst-Case: The quantum term in [26] is formulated as a "worst-case" bound, typically involving a supremum over inputs or hypotheses (e.g., $\sup_{w}\tilde{D}_{\gamma}$). Conversely, our bound is formulated for the average case: it depends on the divergence of the expected classical-quantum state, allowing us to directly incorporate the actual data distribution.
Appendix E Proofs of Theorem 4, Corollary 3 and Corollary 4
E-A Proof of Theorem 4
The proof of Theorem 4 above relies on the Claims 1 and 2 below. The proofs of these claims are given in Appendix J and K.
Claim 1.
Consider $\rho,\rho^{\prime},\sigma\in\mathcal{D}(\mathcal{H}_{A})$. Then, $D(\rho\|\sigma)\leq D(\rho\|\rho^{\prime})+D_{\max}(\rho^{\prime}\|\sigma).$
Claim 2.
Consider $\rho$ and $\sigma$ be two quantum states over Hilbert space $\mathcal{H}$ such that $\rho\ll\sigma$ and $\sigma$ is a finite mixture of probability distributions such that $\sigma=\sum_{b=1}^{m}P(b)\sigma_{b}$, where $\sum_{b=1}^{m}P(b)=1$, and $\rho\ll\sigma_{b}$ for all $b\in[m]$. Then, $D(\rho\|\sigma)\leq\min_{b\in[m]}\left\{D\left(\rho\|\sigma_{b}\right)-\ln P(b)\right\}.$
Since we aim to obtain an upper-bound on $I[S;WB^{\prime}]$, one way to proceed is to use the fact that $I[S;WB^{\prime}]=\min_{\omega^{B}}D(\sigma^{SB}\|\sigma^{S}\otimes\omega^{B})$ (where $B\equiv WB^{\prime}$). Thus,
$I[S;WB^{\prime}]$ $\leq D(\sigma^{SB}\|\sigma^{S}\otimes\omega^{B})$
$=\sum_{s\in\mathcal{S}}P^{\otimes n}_{Z}(s)D(\mathcal{N}^{s}(\rho_{s})\|\omega^{B}).$ (38)
We now choose different values of $\omega^{B}$ to obtain upper-bounds on $I[S;WB^{\prime}]$ discussed in steps below.
(Step $\bf 1$}) Consider $\omega^{B}$ to be a uniform mixture of $\mathcal{N}^{\bf f}(\rho_{\mathbf{f}}),$ over all the types $\mathbf{f}\in T^{n}_{\left|\mathcal{Z}\right|}$ i.e. $\omega^{B}:=\frac{1}{\left|T^{n}_{\left|\mathcal{Z}\right|}\right|}\sum_{\mathbf{f}\in T^{n}_{\left|\mathcal{Z}\right|}}\mathcal{N}^{\bf f}(\rho_{\mathbf{f}})$, then, using Claim 2 and the fact that $\left|T^{n}_{\left|\mathcal{Z}\right|}\right|\leq(n+1)^{\left|\mathcal{Z}\right|-1}$ (see [14, Eq. $6.18$]), Eq. (38) can be upper-bounded as follows,
$~~~I[S;WB^{\prime}]$
$\leq\sum_{s\in\mathcal{S}}P^{\otimes n}_{Z}(s)\min_{\mathbf{f}\in T^{n}_{\left|\mathcal{Z}\right|}}\left\{D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{\bf f}(\rho_{\mathbf{f}}))-\ln\left|T^{n}_{\left|\mathcal{Z}\right|}\right|^{-1}\right\}$
$\leq(\left|\mathcal{Z}\right|-1)\ln(n+1).$
Observe that the above upper-bound on $I[S;WB^{\prime}]$ is independent of the privacy parameters ($\varepsilon,\delta$) of $\mathcal{A}$. This happened because, we chose $\omega^{B}$ to be a uniform mixture of representative quantum states of each type. This choice implied that $\min_{\mathbf{f}\in T^{n}_{\left|\mathcal{Z}\right|}}D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{\bf f}(\rho_{\mathbf{f}}))=0$. To get an upper-bound on $I[S;WB^{\prime}]$ in terms of the privacy parameters, we need to choose $\omega^{B}$ which makes use of the fact that $\mathcal{A}$ satisfies (10). We will accomplish this by using a grid covering for the types of $\mathcal{S}$. We discuss in the step below.
(Step $\bf 2$}) In contrast to Step $1$, we will now choose $\omega^{B}$ to be a mixture over a smaller collection of the output states of $\mathcal{A}$. This smaller collection is obtained by using a grid covering over the types of $\mathcal{S}$, which was developed in the proof of Proposition $2$ of [23]. We now discuss their grid covering over the types of $\mathcal{S}$ below.
Observe that any type ${\bf f}\in T^{n}_{\left|\mathcal{Z}\right|}$ can be thought of as a point inside a $\left|\mathcal{Z}\right|-1$ dimensional grid $0,n]^{\left|\mathcal{Z}\right|-1}$, which is of size $(n+1)^{\left|\mathcal{Z}\right|-1}$. This is because, for any ${\bf f}=({\bf f}_{1},\cdots,{\bf f}_{\left|\mathcal{Z}\right|})\in T^{n}_{\left|\mathcal{Z}\right|}$, the first $\left|\mathcal{Z}\right|-1$ coordinates decide the last coordinate ${\bf f}_{\left|\mathcal{Z}\right|}$, since we have a constraint $\sum_{i=1}^{\left|\mathcal{Z}\right|}{\bf f}_{i}=n$. We now split each dimension of the grid $0,n]^{\left|\mathcal{Z}\right|-1}$ (which is a $[0,n]$ interval) into $t$ equal parts for some
$t\in\mathbb{N}:t\in[1,n].$ (39)
We can think of the grid $0,n]^{\left|\mathcal{Z}\right|-1}$ as a cover of $t^{\left|\mathcal{Z}\right|-1}$ smaller grids of length $l:=\frac{n}{t}$. Note that each side of the smaller grid has $\left\lfloor l\right\rfloor+1$ points. Further, if $\left\lfloor l\right\rfloor+1$ is odd, then we choose the central point of the smaller corresponding to the coordinates of the center of the smaller grid. Thus, for any $s\in\mathcal{S}$, if we consider its type as ${\bf f}^{(s)}$, then we can find a type ${\bf g}^{(s)}\in T^{n}_{\left|\mathcal{Z}\right|}$ such that the first $\left|\mathcal{Z}\right|-1$ coordinates of $\bf f$} are the coordinates of the center of the smaller grid in which the first $\left|\mathcal{Z}\right|-1$ coordinates of ${\bf f}^{(s)}$ resides. In each dimension of the bigger grid, the distance between $s$ and the center of the nearest smaller grid ${\bf c}^{s}$ is given as follows,
$\left|{\bf f}^{(s)}_{z_{(i)}}-{\bf c}^{(s)}_{z_{(i)}}\right|\leq\frac{\left\lfloor l\right\rfloor+1}{2}\leq\frac{n}{2t}+\frac{1}{2},\mbox{~for each }i\in[\left|\mathcal{Z}\right|-1],$
where $z_{(i)}$ is the $i$-th element of the alphabet $\mathcal{Z}$. Therefore, if along all dimension $i\in[\left|\mathcal{Z}\right|-1]$, ${\bf f}^{(s)}_{z_{(i)}}-{\bf g}^{(s)}_{z_{(i)}}=-\frac{n}{2t}+\frac{1}{2}$, then the count of last element $z_{(\left|\mathcal{Z}\right|)}\in\mathcal{Z}$ has to compensate for it. Thus, we have the following,
$\left|{\bf g}^{(s)}_{z_{(\left|\mathcal{Z}\right|)}}-{\bf f}_{z_{(\left|\mathcal{Z}\right|)}}\right|$ $\leq(\left|\mathcal{Z}\right|-1)\left(\frac{n}{2t}+\frac{1}{2}\right).$
Then, for any $s\in\mathcal{S}$ the following holds,
$d(s,T_{{\bf g}^{(s)}})$ $\leq(\left|\mathcal{Z}\right|-1)\frac{n}{t},$ (40)
where $d(s,T_{{\bf g}^{(s)}})$ is distance between the types of $s$ and ${\bf g}^{(s)}$ as defined in Section II.
(Step $\bf 3$}) We now prove Theorem 4 using the grid covering technique discussed in the proof of [23, Proposition 2]. Fix $\omega^{B}=\sum_{{\bf f}\in T^{\prime}}\frac{1}{\left|T^{\prime}\right|}\mathcal{N}^{\bf f}(\rho_{\bf f})$, where $T^{\prime}$ is the collection of the center points of all the smaller grids obtained in Step 2. Then, using Claim 2 and the fact that $\left|T^{\prime}\right|\leq t^{\left|\mathcal{Z}\right|-1}$, we have,
$I[S;WB^{\prime}]$ $\leq\sum_{s\in\mathcal{S}}P^{\otimes n}_{Z}(s)\min_{{\bf f}\in T^{\prime}}\left\{D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{\bf f}(\rho_{\bf f}))+(\left|\mathcal{Z}\right|-1)\ln t\right\}$
$\leq\sum_{s\in\mathcal{S}}P^{\otimes n}_{Z}(s)\left(D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{{\bf g}^{(s)}}(\rho_{{\bf g}^{(s)}}))+(\left|\mathcal{Z}\right|-1)\ln t\right).$ (41)
(Step $\bf 4$}) We will now analyze the first term in the RHS of (41) by using Claim 1 and [24, Lemma $6.9$]. Toward this, in [24, Lemma $6.9$], let $\rho=\mathcal{N}^{s}(\rho_{s})$ and $\sigma=\mathcal{N}^{{\bf g}^{(s)}}(\rho_{{\bf g}^{(s)}})$. Thus, [24, Lemma $6.9$] implies that there exists a quantum state $\mathcal{N}^{s}(\rho_{s})^{\prime}$ in the close vicinity of $\mathcal{N}^{s}(\rho_{s})$ such that $D_{\max}(\mathcal{N}^{s}(\rho_{s})^{\prime}\|\mathcal{N}^{{\bf g}^{(s)}}(\rho_{{\bf g}^{(s)}}))\leq f(\varepsilon,\delta)$, where $f(\cdot,\cdot)$ is some function. Thus, using [24, Lemma $6.9$], Claim 1, Assumption (12) and the extension of privacy constraints of $\mathcal{A}$ under $k$-neighboring inputs, we have the following series of inequalities,
$D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{{\bf g}^{(s)}}(\rho_{{\bf g}^{(s)}}))$ $\leq D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{s}(\rho_{s})^{\prime})+\varepsilon^{\prime}$
$\leq\frac{2}{m}E^{2}_{1}\left(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{s}(\rho_{s})^{\prime}\right)+\varepsilon^{\prime}$
$\leq\varepsilon^{\prime}+\frac{2}{m}g_{\frac{n(\left|\mathcal{Z}\right|-1)}{t}}(\varepsilon,\delta),$ (42)
where $\varepsilon^{\prime}:=\frac{n(\left|\mathcal{Z}\right|-1)\varepsilon}{t}+\ln\frac{1}{1-g_{\frac{n(\left|\mathcal{Z}\right|-1)}{t}}(\varepsilon,\delta)}$, and $g_{\frac{n(\left|\mathcal{Z}\right|-1)}{t}}(\varepsilon,\delta)=\frac{e^{\frac{n(\left|\mathcal{Z}\right|-1)\varepsilon}{t}}-1}{e^{\varepsilon}-1}\delta$. Thus, using Eqs. (41) and (42) we have,
$\hskip-7.0ptI[S;WB^{\prime}]{\leq}\frac{n(\left|\mathcal{Z}\right|-1)\varepsilon}{t}+(\left|\mathcal{Z}\right|-1)\ln t+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta),$ (43)
where $h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta):=\ln\frac{1}{1-g_{{n(\left|\mathcal{Z}\right|-1)}}(\delta)}+\frac{2}{m}g_{n(\left|\mathcal{Z}\right|-1)}(\delta)$ (observe that $h_{\left|\mathcal{Z}\right|}(\varepsilon,0)=0$) and the last inequality follows from the fact that the grid size $t\geq 1$.
(Step $\bf 5$}) In this step, we optimize the choice over $t$ (grid size) to tighten the upper-bound obtained in Eq. (43). Observe that the value of $t$ which minimizes the RHS of Eq. (43) is,
$t^{\star}=n\varepsilon.$ (44)
As mentioned in the statement of Theorem 4, we have $\varepsilon\in\left[\frac{1}{n}\right.,1]$ and thus (44) yields that $1\leq t^{\star}\leq n$, which satisfies the size constraint of grid mentioned in (39). Therefore, substituting $t=t^{\star}$ in (43) yields,
$I[S;WB^{\prime}]$ $\leq(|\mathcal{Z}|-1)\left(1+\ln\left({n\varepsilon}\right)\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)$
$=(|\mathcal{Z}|-1)\ln\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$
This completes the proof of Theorem 4.
E-B Proof of Corollary 3
For $\varepsilon<\frac{1}{n}$ (as mentioned in Corollary 3), (44) yields that $t^{\star}<1$, and therefore it does not (39). Thus, in this case, we set $t=1$ in (43) to obtain the desired upper-bound.
E-C Proof of Corollary 4
For $\varepsilon>1$ (as mentioned in Corollary 4), (44) yields that $t^{\star}>n$, which does not satisfy the grid size constraint mentioned in (39). Therefore in this case we set grid size $t=n$. However, for this choice of grid size, observe that the grid covers all the sequences in $\mathcal{S}$ and therefore covers all the type-representatives in $\mathcal{S}$. This is the same case as Step 1. Therefore, we have,
$I[S;WB^{\prime}]$ $\leq(\left|\mathcal{Z}\right|-1)\ln(n+1).$
This completes the proof of Corollary 4. Further, note that if we substitute $t=n$ in (43), then it would yield us a weaker bound as compared to the above.
Appendix F Comparison of Upper-bounds on Stability with Prior Work
In this section, we compare the stability upper-bound (Theorem 4) derived in this work with existing results.
F-A Comparison between Theorem 4 and [20, Proposition 10]
In [20, Proposition 10], the authors derived an upper bound on the Holevo information for quantum $(\varepsilon,\delta)$-LDP quantum channels, as stated in Eq. (209) of [20]. However, one of the authors of [20] later clarified to the authors of the present paper [28] that the phrase “for quantum $(\varepsilon,\delta)$-LDP quantum channels” was a typographical error.
The corrected statement is as follows. If an algorithm $\mathcal{A}$ satisfies $\varepsilon$-QLDP, meaning that
$\mathrm{Tr}[M\mathcal{A}(\rho_{x})]\leq e^{\varepsilon}\mathrm{Tr}[M\mathcal{A}(\rho_{x^{\prime}})],\qquad\forall\,x,x^{\prime}\in\mathcal{X},\ \forall\,M:0\leq M\leq I,$ (45)
then, the following bound holds:
$I[X;B]_{\sigma}\leq\varepsilon\tanh\left(\frac{\varepsilon}{2}\right)=\varepsilon\left(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\right),$ (46)
where $I[X;B]_{\sigma}$ is the Holevo information computed with respect to the state $\sigma:=\sum_{x\in\mathcal{X}}P(x)|x\rangle\langle x|\otimes\mathcal{A}(\rho_{x})$. Here, their learning algorithm $\mathcal{A}$ can be considered as a map $\mathcal{N}$ given in (3), but the map does not depend on $s$. Hence, the reference [20] also adopts the Untrusted Data Processor scenario similar to the second part of [8].
In contrast, our main result, Theorem 4, provides an upper bound under a weaker assumption: the algorithm $\mathcal{A}$ satisfies $1$-neighbor $(\varepsilon,\delta)$-DP, i.e., the condition (10) holds for every $s\overset{1}{\sim}s^{\prime}$ (see Section II). When we set $\delta=0$ and $\mathcal{X}=\mathcal{S}$, the form of the constraint in (10) becomes identical to that in (45). However, our result applies this constraint only to $1$-neighboring pairs, whereas their result assumes it for all pairs $s\neq s^{\prime}\in\mathcal{S}$.
In fact, if we strengthen our assumption to match theirs—namely, require (10) for any distinct $s,s^{\prime}\in\mathcal{S}$—then part $(i)$ of [9, Corollary 3] recovers the same bound as (46), thereby aligning our result with the corrected version of [20, Proposition 10].
The proof of Theorem 1 formally establishes the connection between algorithmic stability and generalizability by treating the mutual information $I[S;WB^{\prime}]$ as a proxy for stability. We first demonstrate that the expected generalization error is fundamentally limited by the square root of the information the algorithm leaks about the training data $S$, i.e., we have the following,
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ $\leq\sqrt{2\alpha^{2}I[S;WB^{\prime}]}.$ (47)
Here, $I[S;WB^{\prime}]$ quantifies the dependence of the output hypothesis on the specific training set; a lower value implies that the algorithm is "stable" and not overfitting to individual data points. The crucial link to Theorem 4 is that it provides the explicit upper bound on this stability measure derived solely from the privacy constraints. Finally, by substituting the bound from Theorem 4 into (47), we mathematically confirm that the rigorous stability imposed by $(\varepsilon,\delta)$-DP directly suppresses the generalization error, by preventing the algorithm from depending too heavily on any single data point and ensuring that the learned hypothesis performs well on unseen data.
F-B Comparison between Theorem 4 and [8, Appendix C.7]
The reference [8] studies the local differential privacy of learning algorithms in two settings. That is, their discussion is composed of two parts, the first part starting with “First” and the second part starting with “Next”.
F-B1 Assumption in [8, Appendix C.7]
Their first part discusses the Holevo information under a certain condition. However, a careful examination of the proof in [8, Appendix C.7] reveals that the argument relies on a stronger assumption than their statement as follows. In this place, the authors claim to prove the following bound on the Holevo information.
$I(\text{test};\text{hyp})_{\sigma^{\mathcal{A}}_{(s,w)}}\leq 2\varepsilon(1-e^{-\varepsilon})\sqrt{2I(\text{test};\text{train})_{\rho^{\mathcal{A}}_{(s,w)}}},$ (48)
under the assumption that the channel $\Lambda^{\mathcal{A}}_{s,w}:\mathcal{H}^{\mbox{train}}\to\mathcal{H}^{\mbox{hyp}}$ is $\varepsilon$-LDP, i.e.,
$\mathrm{Tr}\left[M\Lambda^{\mathcal{A}}_{s,w}(\rho^{\mbox{train}}_{1})\right]\leq e^{\varepsilon}\mathrm{Tr}\left[M\Lambda^{\mathcal{A}}_{s,w}(\rho^{\mbox{train}}_{2})\right],$ (49)
for all $0\preceq M\preceq\mathbb{I}^{\mbox{hyp}}$ and $\rho^{\mbox{train}}_{1},\rho^{\mbox{train}}_{2}\in\mathcal{D}(\mathcal{H}^{\mbox{train}})$.
However, a closer examination of their proof reveals that the argument implicitly depends on a stronger condition, namely
$\mathrm{Tr}\left[O\big(\mathbb{I}^{\mathrm{test}}\otimes\Lambda^{\mathcal{A}}_{s,w}\big)\big(\rho^{\mathrm{test;train}}_{1}\big)\right]\leq e^{\varepsilon},\mathrm{Tr}\left[O\big(\mathbb{I}^{\mathrm{test}}\otimes\Lambda^{\mathcal{A}}_{s,w}\big)\big(\rho^{\mathrm{test;train}}_{2}\big)\right],$ (50)
for all $0\preceq O\preceq\mathbb{I}^{\mathrm{test;hyp}}$ and $\rho^{\mathrm{test;train}}_{1},\rho^{\mathrm{test;train}}_{2}\in\mathcal{D}(\mathcal{H}^{\mathrm{test;train}})$. In other words, the proof appears to require that $\Lambda^{\mathcal{A}}_{s,w}$, which is locally $\varepsilon$-LDP on $\mathcal{H}^{\mathrm{hyp}}$, also preserves $\varepsilon$-LDP globally when extended to the joint space $\mathcal{H}^{\mathrm{test;hyp}}$. Crucially, (49) does not imply (50). Indeed, by [30, Theorem 4], the identity channel on $\mathcal{H}^{\mathrm{test}}$ fails to satisfy differential privacy for any $\varepsilon\geq 0$, so the composition $\mathbb{I}^{\mathrm{test}}\otimes\Lambda^{\mathcal{A}}_{s,w}$ cannot satisfy $\varepsilon$-LDP solely on the basis of (49). Therefore, there is a gap in the argument of [8, Appendix C.7]: the claimed bound (48) does not follow from their stated assumption (49). That is, one must assume (50) instead of (49). Moreover, the upper bound obtained in (48) involves the term $I(\text{test};\text{train})_{\rho^{\mathcal{A}}{(s,w)}}$. To render this bound meaningful, $I(\text{test};\text{train})_{\rho^{\mathcal{A}}{(s,w)}}$ should also be controlled by some function of the security parameter $\varepsilon$, although our evaluations—such as Theorem 4—do satisfy this requirement.
F-B2 Security condition in [8, Appendix C.7]
Their second part essentially changes their model into the Untrusted Data Processor scenario studied in Section V because on the page $59$ of [8] the authors mention the following:
“Next, we turn our attention to the classical MI term in our generalization bounds. Here, we assume that the learner $\mathcal{A}$ uses an overall $\varepsilon$-LDP POVM. As the POVM $\left\{|s\rangle\langle s|\otimes E_{s}^{\mathcal{A}}(w)\right\}_{s,w}$ is not LDP even if every $\left\{E_{s}^{\mathcal{A}}(w)\right\}_{w}$ is, we make the simplifying assumption that the learner uses an $s$-independent $\varepsilon$-LDP POVM $\left\{E^{\mathcal{A}}(w)\right\}_{w}$.”
Even in this scenario, our results still hold, as explained in Section V. However, in this scenario, it is reasonable to impose the ITA condition given in Definition 7 to our learning algorithm, as discussed in Section V while they did not consider such a constraint.
Appendix G Technical Discussion related to Information-Theoretic Admissibility (ITA)
In the scenario where the Data Processor cannot be trusted, the privacy guarantee must hold against the Data Processor itself. Unlike the trusted setting where the Investigator only sees the classical output $w$, here the adversary has access to the quantum output $\mathcal{N}(\rho_{s})$. Consequently, the definition of differential privacy must be adapted to constrain the indistinguishability of the quantum states output by the learning map directly.
Definition 11.
An algorithm $\mathcal{N}$ is said to be a $1$-neighbor $(\varepsilon,\delta)$-DP support-consistent learning algorithm with an untrusted Data Processor if the following conditions hold.
  1. 1.
    Permutation Invariance: For all $s,s^{\prime}\in\mathcal{S}$ satisfying $T_{s}=T_{s^{\prime}}$, the algorithm satisfies the condition $\mathcal{N}(\rho_{s})=\mathcal{N}(\rho_{s^{\prime}})$.
  2. 2.
    Privacy: For every $s\overset{1}{\sim}s^{\prime}$ (see Section II) and $0\preceq\Lambda\preceq\mathbb{I}$, the following inequality holds:
    $\mathrm{Tr}[\Lambda\mathcal{N}(\rho_{s})]$ $\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}(\rho_{s^{\prime}})]+\delta,$
    $\mathrm{Tr}[\Lambda\mathcal{N}(\rho_{s^{\prime}})]$ $\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}(\rho_{s})]+\delta.$
  3. 3.
    Support Consistency: For every $s\overset{1}{\sim}s^{\prime}$, the output supports are identical, i.e.,
    $\operatorname{supp}(\mathcal{N}(\rho_{s}))=\operatorname{supp}(\mathcal{N}(\rho_{s^{\prime}})).$ (51)
To rigorously assess whether the privacy constraints are meaningful, we must ensure that the learning algorithm is not artificially “noisy” or suboptimal. We formalize this via the notion of admissibility. This concept allows us to order channels by their information content and identify when an algorithm extracts all accessible information from the input states.
Definition 12 (Information ordering of Algorithms).
Let $\mathcal{N}:=\{\mathcal{N}_{w}\}_{w}$ and $\mathcal{N}^{\prime}:=\{\mathcal{N}^{\prime}_{w}\}_{w}$ be two quantum learning algorithms and $\{\rho_{s}\}_{s}$ be a fixed set of input states to these algorithms. We say that $\mathcal{N}^{\prime}$ is more informative than $\mathcal{N}$ with respect to $\{\rho_{s}\}_{s}$ if there exists a family of CP-TP maps $\{\Gamma_{w}\}_{w}$ such that
$\Gamma_{w}\circ\mathcal{N}^{\prime}_{w}(\rho_{s})=\mathcal{N}_{w}(\rho_{s}),\quad\forall\,(s,w)\in\mathcal{S}\times\mathcal{W}.$
Furthermore, $\mathcal{N}^{\prime}$ is said to be strictly more informative than $\mathcal{N}$ with respect to $\{\rho_{s}\}_{s}$ if $\mathcal{N}^{\prime}$ is more informative than $\mathcal{N}$, but the converse does not hold.
G-A Implications of ITA: Source-Layer Privacy and Quantum Advantage
Resolving the ITA Conflict: Source-Layer Privacy. The impossibility result in Lemma 1 implies that in the classical domain, if a Data Processor is untrusted and executes an ITA (optimal) algorithm, privacy cannot be preserved by the algorithm itself. Since an ITA algorithm extracts all available information, the output effectively reveals the raw input. Consequently, to preserve privacy in the classical untrusted setting, the burden of protection must shift from the algorithm to the input data itself. This is standardly achieved via Input Perturbation or Local Differential Privacy (LDP), where the Respondent applies a local randomization mechanism $\mathcal{M}$ to generate a noisy version $\tilde{s}=\mathcal{M}(s)$. Even if the Untrusted Processor fully recovers $\tilde{s}$ (as allowed under ITA), the underlying sensitive data $s$ remains protected by the noise added at the source. Thus, indistinguishability is enforced at the source layer, making the specific choice of the processor’s algorithm irrelevant to the privacy guarantee.
Quantum Encoding as Intrinsic Source Noise. This necessity for source-layer protection provides a rigorous motivation for our quantum learning framework. In our model, the encoding map $s\mapsto\rho_{s}$ plays a role conceptually equivalent to classical input perturbation, but with a fundamental physical advantage. In the classical setting, distinct data points $s\neq s^{\prime}$ are perfectly distinguishable unless artificial noise is added. In the quantum setting, however, if the encoded states $\{\rho_{s}\}_{s}$ are non-orthogonal, they are physically indistinguishable with certainty. This non-orthogonality introduces an intrinsic, unavoidable uncertainty—effectively “quantum noise”—that prevents even an adversary with unlimited computational power from perfectly distinguishing $s$ from $s^{\prime}$. Therefore, our framework intrinsically embeds privacy into the physical layer. Even if the Untrusted Data Processor employs an ITA algorithm (i.e., performs the optimal Helstrom measurement to extract maximum information), their ability to infer $s$ is fundamentally limited by the non-orthogonality of the encoded states. This confirms that our security condition is robust: privacy is not contingent on the Processor’s cooperation but is guaranteed by the physical nature of the encoding itself.
Appendix H Proof of Corollary 2
Since $s\overset{k}{\sim}s^{\prime}$, there exists a $k+1$-length sequence $\{s_{i}\}_{i=0}^{k}\subseteq\mathcal{S}$ such that $s_{0}=s$, $s_{k}=s^{\prime}$ and for each $i\in[k]$, $s_{i-1}\overset{1}{\sim}s_{i}$. Thus, for any $0\preceq\Lambda\preceq\mathbb{I}$, using Eq (10), we have
$\mathrm{Tr}[\Lambda\mathcal{N}^{(s)}(\rho_{s})]$ $\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s_{1})}(\rho_{s_{1}})]+\delta$
$\leq e^{2\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s_{2})}(\rho_{s_{2}})]+(e^{\varepsilon}+1)\delta$
$\leq e^{3\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s_{3})}(\rho_{s_{3}})]+(e^{2\varepsilon}+e^{\varepsilon}+1)\delta$
[…]
$\leq e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}^{(s^{\prime})}(\rho_{s^{\prime}})]+(e^{(k-1)\varepsilon}+e^{(k-2)\varepsilon}+\cdots+e^{\varepsilon}+1)\delta$
$=e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}(\sigma)]+g_{k}(\delta).$
This completes the proof of Corollary 2.
Appendix I Proof of Lemma 1
Choose a basis $\{|x\rangle\}$ that diagonalizes $\rho_{s}$. In this basis, the state can be expressed as
$\rho_{s}=\sum_{x}P_{X|s}(x)\,|x\rangle\langle x|,$ (52)
on system $X$. Next, define the instrument $\{\mathcal{N}_{w}^{\prime}\}_{w}$ by
$\mathcal{N}_{w}^{\prime}(|x\rangle\langle x|):=|x\rangle\langle x|\otimes\mathcal{N}_{w}(|x\rangle\langle x|),$ (53)
where the output system is $XB^{\prime}$. Recall that the original learning algorithm $\{\mathcal{N}_{w}\}_{w}$ outputs on system $B^{\prime}$. Under this construction, $\{\mathcal{N}_{w}^{\prime}\}_{w}$ is strictly more informative than $\{\mathcal{N}_{w}\}_{w}$.
Suppose, for contradiction, that $\{\mathcal{N}_{w}\}_{w}$ is more informative than $\{\mathcal{N}_{w}^{\prime}\}_{w}$. Then there exist CP-TP maps $\{\Gamma_{w}\}_{w}$ such that
$\Gamma_{w}(\mathcal{N}_{w}(\rho_{s}))=\mathcal{N}_{w}^{\prime}(\rho_{s})\quad\text{for all }s\in\mathcal{S}.$
Hence,
$\mathrm{Tr}_{B^{\prime}W}\Bigg[\sum_{w}\Gamma_{w}(\mathcal{N}_{w}(\rho_{s}))\otimes|w\rangle\langle w|\Bigg]$
$=$ $\mathrm{Tr}_{B^{\prime}W}\Bigg[\sum_{w}\mathcal{N}_{w}^{\prime}\Bigg(\sum_{x}P_{X|s}(x)|x\rangle\langle x|\Bigg)\otimes|w\rangle\langle w|\Bigg]$
$=$ $\mathrm{Tr}_{B^{\prime}W}\Bigg[\sum_{x}P_{X|s}(x)\sum_{w}\mathcal{N}_{w}^{\prime}(|x\rangle\langle x|)\otimes|w\rangle\langle w|\Bigg]$
$=$ $\mathrm{Tr}_{B^{\prime}W}\Bigg[\sum_{x}P_{X|s}(x)|x\rangle\langle x|\otimes\sum_{w}\mathcal{N}_{w}(|x\rangle\langle x|)\otimes|w\rangle\langle w|\Bigg]$
$=$ $\sum_{x}P_{X|s}(x)|x\rangle\langle x|=\rho_{s},$ (54)
which contradicts the assumption that no CP-TP map $\Gamma$ satisfies
$\Gamma\left(\sum_{w\in\mathcal{W}}\mathcal{N}_{w}(\rho_{s})\otimes|w\rangle\langle w|\right)=\rho_{s}.$
Therefore, $\{\mathcal{N}_{w}^{\prime}\}_{w}$ is strictly more informative than $\{\mathcal{N}_{w}\}_{w}$, and thus $\{\mathcal{N}_{w}\}_{w}$ is not ITA.
Appendix J Proof of Claim 1
By the definition of the max-relative entropy, we have
$\rho^{\prime}\leq e^{D_{\max}(\rho^{\prime}\|\sigma)}\,\sigma.$
Equivalently,
$\sigma\geq e^{-D_{\max}(\rho^{\prime}\|\sigma)}\,\rho^{\prime}.$
Since the logarithm is operator monotone, this implies
$\ln\sigma\succeq\ln\rho^{\prime}-D_{\max}(\rho^{\prime}\|\sigma)\,\mathbb{I}.$
Multiplying both sides by $\rho$ and taking the trace (which reverses the inequality), we obtain
$\mathrm{Tr}[\rho\ln\sigma]\leq-\mathrm{Tr}[\rho\ln\rho^{\prime}]+D_{\max}(\rho^{\prime}\|\sigma).$
Adding $\mathrm{Tr}[\rho\ln\rho]$ to both sides gives
$\mathrm{Tr}[\rho(\ln\rho-\ln\sigma)]\leq\mathrm{Tr}[\rho(\ln\rho-\ln\rho^{\prime})]+D_{\max}(\rho^{\prime}\|\sigma),$
which can be written as
$D(\rho\|\sigma)\leq D(\rho\|\rho^{\prime})+D_{\max}(\rho^{\prime}\|\sigma).$
This completes the proof of Claim 1.
Appendix K Proof of Claim 2
We begin by invoking the operator monotonicity of the function $\ln(\cdot)$. Since $\sigma\succeq P(b)\sigma_{b}$ for every $b$, we obtain
$\ln\sigma\succeq\ln P(b)\,\mathbb{I}+\ln\sigma_{b}.$ (55)
Using (55), we immediately have,
$D(\rho\|\sigma)$ $=\mathrm{Tr}\left[\rho(\ln\rho-\ln\sigma)\right]$
$\leq\mathrm{Tr}\left[\rho(\ln\rho-\ln P(b)\,\mathbb{I}-\ln\sigma_{b})\right]$
$=D(\rho\|\sigma_{b})-\ln P(b).$ (56)
Since (56) holds for all $b$, it implies Claim 2.